### v1

# Large Language Models Can Learn Temporal Reasoning

Abstract
Large language models (LLMs) learn temporal concepts from the co-occurrence of related tokens in a sequence. Compared with conventional text generation, temporal reasoning, which reaches a conclusion based on mathematical, logical and commonsense knowledge, is more challenging. In this paper, we propose TempGraph-LLM, a new paradigm towards text-based temporal reasoning. To be specific, we first teach LLMs to translate the context into a temporal graph. A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for pre-training on this task. We prove in experiments that LLMs benefit from the pre-training on other tasks. On top of that, we guide LLMs to perform symbolic reasoning with the strategies of Chain of Thoughts (CoTs) bootstrapping and special data augmentation. We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free text.
1 Introduction
As one of the fundamental abilities, temporal reasoning plays an important role in human perception. It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023). However, they still can not perform temporal reasoning very well as we expected  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023). Compared with other singular types, temporal reasoning is a form of integrated task built upon foundations like mathematics, logical and commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
Figure 1: Our TempGraph-LLM framework has two steps: text-to-tempgraph translation and neuro-symbolic reasoning.
Recent work mainly adopts the general methods to investigate and improve the temporal reasoning capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different temporal reasoning tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include span extraction pre-training, instruction tuning and reinforcement learning.
Despite the effectiveness of such methods, they ignore the intrinsic nature of temporal reasoning, and cannot guarantee the correctness of intermediate steps  Saparov and He (2023); Yang et al. (2023b). Human performs temporal reasoning on a timeline of events which are aligned with the entities and relations. These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information. Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline from the text, and then to perform deliberate reasoning on it. However, the high-quality extraction of aligned timeline is a challenging task itself due to lack of ground truth. To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision. We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks.
Given the reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on that. (1) How to introduce the necessary arithmetic and commonsense knowledge involved in temporal reasoning? (2) How to guide LLMs to perform rigorous reasoning? Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap of the timestamps, the relative order of the time gaps). As for (2), there are always two roadmaps: (i) translating natural language into logical statements and using symbolic inference engine for reasoning  Pan et al. (2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the generated text especially when training data are insufficient  (Yang et al., 2023b). To solve these problems, we propose a symbolic integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. Furthermore, we design two types of data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
To be specific, our contributions are summarized as follows:
  • We propose a new paradigm, TempGraph-LLM, for text-based temporal reasoning. In this framework, we first translate the context into a temporal graph, and then perform symbolic reasoning on it. Extensive experiments prove that it brings better performance than convetional methods.
  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLMs to utilize symbolic reasoning, which brings better performance than free text.
  • We present a pipeline to create a synthetic dataet for question answering that requires temporal reasoning. It is fully controllable and requires minimal supervision for context-tempgraph alignment. We show in experiments that pre-training on our dataset benefits LLMs on other tasks.
2 Dataset Construction
We propose a pipeline for the construction of our synthetic temporal QA dataset called TempGraph QA. Each sample, as shown in Table 1, is in the form of (story, tempgraph, question, answer). The tempgraph is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018). The number of events in each tempgraph is restricted into a certain range. By using some rule-based templates, we obtain reliable question and answer pairs from the graph. Finally, each story is translated from the tempgraph by GPT-3.5. Below we show the details for all steps.
TempGraph (sub; rel; obj; start/end; time):
[…]
[2] (John Thompson owned Pearl Network) starts at 1942;
[3] (Sophia Parker was married to John Thompson) starts
at 1947; [4] (John Thompson was married to Sophia Parker)
starts at 1947; [5] (Sophia Parker was married to John
Thompson) ends at 1953; [6] (John Thompson was married
to Sophia Parker) ends at 1953; [7] (John Thompson owned
Pearl Network) ends at 1967; [8] (John Thompson died in
Riverside) starts at 1988; [9] (Sophia Parker died in
Lancaster) starts at 1995.
Graph-based Story (from GPT-3.5):
[…]
boy named John Thompson was brought into the world
in the year 1921. Growing up, he had a vibrant spirit
and an adventurous soul. $\cdots$
Graph-based QA (from Rule-based Python Script):
[…]
Pearl Network) or (John Thompson was married to
Sophia Parker)?
A: (John Thompson owned Pearl Network).
Q: True or false: event (John Thompson owned Pearl
Network) was longer in duration than event (Sophia
Parker was married to John Thompson)?
A: True.
$\cdots$
Table 1: Each sample from TempGraph QA is in the form of (story, tempgraph, question, answer).
Reasoning Type Question Answer
Sequencing Which event occurred first, {event A} or {event B}? {event A}/{event B}
Given the following five events: {event A}, {event B}, {event C}, {event D}, {event A}/{event B}/
{event E}, which event is the first/second/third/fourth/fifth one in chronolog- {event C}/{event D}/
ical order? {event E}
Duration How long did the {event A} last? {duration of event A}
Temporal How much time passed between the start of {event A} and the {gap between event A
Relation start of {event B}? and B’s start time}
What happened right before/after {event A} started? {event B} right before/
after {event A}
Facts When did the {event A} occur? {event A’s start time}
Extraction
Simultaneity True or false: {event A} and {event B} happened at the same year? True / False
True or false: {event A} was still happening when {event B} started? True / False
Comparative True or false: {event A} was longer in duration than {event B}? True / False
Analysis
Table 2: Reasoning types with corresponding questions and answers in TempGraph QA.
Step 1: Graph Splitting & Masking   Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them all at once. Given a specific entity, we extract its third-order neighbors and all the events happening between them. To avoid the discrepancy caused by the number of events, we set a certain range and discard the non-compliant subgraphs. Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training and test samples. We finally obtain 500 tempgraphs for training and another 100 for test. In addition, we consider the noticeable leakage problem of LLMs, where the prior knowledge of test data has been implicitly obtained from the extensive pre-training. Thus, a masking strategy, i.e., changing entities names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5. To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
Step 2: Graph-based Open QA Creation   Based on the given tempgraph, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we split the starting and ending of a persistent event into two transient ones. To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design nine different question types (shown in Table 2), and write a rule-based Python script to determine the answers from the tempgraph.
Step 3: Quality Control   Although LLMs are good at story generation, there is a risk of misalignment between the given tempgraph and the generation text. To avoid the interference caused by misalignment, we propose an automatic way for verification. Given a generated story, we ask LLMs the time of each event in the graph. If LLMs cannot give the correct answer, we remove the possibly missing events in the graph. Since this strategy tends to simplify the reasoning, we manually verify the existence of the possibly missing events for both training and test sets. Note that only the quality of story-tempgraph alignment needs control, since all the QAs are generated from the tempgraph with specific rules. We show dataset statistcis in Appendix A, and all the prompts involved in Appendix C.
3 TempGraph-LLM
Motivated by human perception, we propose a new paradigm called TempGraph-LLM. We first translate the text into a temporal graph, and then guide LLMs to perform symbolic reasoning on it.
3.1 Text-to-TempGraph Translation
TempGraph (TG) serves as the foundation of the following symbolic reasoning process. Although LLMs can extract them from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs. Thus, we provide a deliberate pipeline for high-quality TG translation.
3.1.1 Pre-training
In the TempGraph QA, each story is corresponding to a verified TG. Similarly, for other data, where the ground truth of TG is avaiable, we can pre-train LLMs on them. Since language models (LMs) can only deal with sequence of tokens, we sort all the events in the graph according to their time. In addition, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events. With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023b) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the pre-training on our data for other tasks.
3.1.2 Fine-tuning
Generally, the biggest challenge for fine-tuning on real-world datasets is lack of ground truth. Instead of directly using the estimation from LLMs, we provide a pipeline for better alignment.
Entity & Relation Extraction   To align the entities and relations in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context. To obtain the entities and relations in pre-defined QAs, we consider two ways: parsing the QAs with rules or using LLMs. The former is more efficient and reliable when the QAs are generated by templates  (Chen et al., 2021; Tan et al., 2023a), while the latter should be used when the QAs are in free text  (Zhou et al., 2019; Ning et al., 2020).
Temporal Info Identification   Compared with conventional reasoning tasks, temporal information is one of the most important and unique components here. To identify all the time expressions, LLMs are first used to extract them in the context, and then a rule-based Python script  (Yang et al., 2023a) is adopted for time-normalization and invalid output filtering. For example, we convert "Dec. 5th, 2004" into "2004-12-05", "July 1989" into "1989-07-XX", and "1973" into "1973-XX-XX", respectively. For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, e.g., "1974 is before 1983", "the time gap between May 2004 and Oct 2008 is 4 years and 5 months", which are introduced in temporal reasoning as external knowledge.
Graph Construction & Verification   We generate the graph using LLMs with in-context learning (ICL). To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output. The TG should only contain the extracted elements to guarantee the alignment with pre-defined QAs. We classify the stories according to whether they contain explicit expressions such as "July 1989", "2003–2005", "the late 1900s". For those without explicit expressions, we ask LLMs to provide the temporal order of different events. We further automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph. For those graphs where LLMs give the wrong answers, we manually check and correct them.
3.2 Neuro-symbolic Temporal Reasoning
Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by Chain of Thoughts (CoTs) bootstrapping and data augmentation.
3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFTs on reliable CoTs bring better performance in reasoning tasks than those on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). However, it is challenging to obtain reliable CoTs from LLMs for symbolic reasoning tasks  (Saparov and He, 2023). Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples. We bootstrap the generations and only accept the ones with correct final answers.
To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{m}\}$ with LLMs. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask LLMs to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$. These CoTs will be manually verified and used as few-shot ICL examples. When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c^{k}_{j^{\prime}},\hat{a}^{k}_{j^{\prime}})$ from LLMs, and only accept the one leading to correct answers, i.e., $\hat{a}^{k}_{j^{\prime}}=a^{*}_{j^{\prime}}$. Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method is more flexible and robust. In addition, inspired by  (Wang et al., 2023), we compare different type of CoTs using contrastive learning. The core idea is to choose CoTs that make the correct answer more plausible than other candidates.
$c^{*}=\operatorname{argmax}\log P\left(a^{*}\mid g,e,q,c\right)+\gamma G\left(c\right)$ (1)
$G\left(c\right)=\log\frac{P\left(a^{*}\mid g,e,q,c\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c\right)}$ (2)
$P\left(a\mid g,e,q,c\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c,t_{<l}\right)$ (3)
where $g,e,q,c$ denote the TG, external knowledge, question, and CoTs, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
3.2.2 Data Augmentation
Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023). To address the issue, we propose two special strategies for data augmentation.
Learning with Disturbances   Essentially, our framework involves two steps: text-to-tempgraph translation and symbolic temporal reasoning. At the test stage, reasoning happens on the noisy graphs, which can impact the overall performance. Specifically, for the reasoning module, we train with ground truth/verified graphs while testing on the estimations. To improve the robustness of reasoning, we introduce some disturbances on the graphs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the model. We investigate two types of disturbances: remove irrelevant edges or replace edges by changing relation names, e.g., "owned" $\rightarrow$ "run", "was affiliated to" $\rightarrow$ "belonged to", "played for" $\rightarrow$ "joined". An edge is considered as irrelevant if it is not involved in both QA and CoTs.
$F_{irr}\in\{F:P(c|g_{\setminus\{F\}},e,q)\approx P_{0}\}_{F\in g}$ (4)
$F^{\prime}\in\{f_{R}(F):P(c|f_{R}(g),e,q)\approx P_{0}\}_{F\in g}$ (5)
TempGraph (Adding Disturbances):
[…]
[2] (John Thompson run Pearl Network) starts at 1942;
[3] (Sophia Parker and John Thompson became life
partner) starts at 1947; [4] (John Thompson and Sophia
Parker became life partner) starts at 1947; [5] (Sophia
Parker and John Thompson became life partner)
ends at 1953; [6] (John Thompson and Sophia Parker)
became life partner) ends at 1953; [7] (John Thompson
run Pearl Network) ends at 1967.
Graph-based QA (No Change):
[…]
Network) was longer in duration than event (Sophia
Parker was married to John Thompson)?
CoT (symbolic): The duration for each event can be
calculated as follows:
(John Thompson owned Pearl Network) starts at
1942, ends at 1967, 1967 - 1942 = 25
(Sophia Parker was married to John Thompson) starts
at 1947, ends at 1953, 1953 - 1947 = 6
25 is greater than 6 , thus, the answer is True.
A: True.
TempGraph (Inductive Transformation):
[…]
[2] (James Brown owned Iris Inn) starts at 1952;
[3] (Ella Perry was married to James Brown) starts
at 1957; [4] (James Brown was married to Ella Perry)
starts at 1957; [5] (Ella Perry was married to James
Brown) ends at 1963; [6] (James Brown was married
to Ella Perry) ends at 1963; [7] (James Brown owned
Iris Inn) ends at 1977; [8] (James Brown died in
Auckland) starts at 1998; [9] (Ella Perry died in
Monaco) starts at 2005.
Graph-based QA (Inductive Transformation):
[…]
was longer in duration than event (Ella Perry was
married to James Brown)?
CoT (symbolic): The duration for each event can be
calculated as follows:
(James Brown owned Iris Inn) starts at 1952, ends at
1977, 1977 - 1952 = 25
(Ella Perry was married to James Brown) starts at
1957 , ends at 1963, 1963 - 1957 = 6
25 is greater than 6 , thus, the answer is True.
A: True.
Table 3: We propose two special data augmentation strategies (adding disturbances and inductive transformation) for the neuro-symbolic reasoning.
where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation name changing $f_{R}(\cdot)$.
Inductive Transformation  In symbolic reasoning, a rule is considered as inductive if it is generalizable to different entities and times  (Xiong et al., 2022, 2023). For LLMs, this might be difficult to understand without enough training data since their fundamental principle is to learn the co-occurrence of tokens. To solve this problem, we introduce a new approach called inductive transformation. We globally map all the entity names in training data to some random names of the same type (e.g., "Knox Cunningham" $\rightarrow$ "Olivia Davis"; "Sydney" $\rightarrow$ "Cardiff"), and change the times based on a global time offset (e.g., the offset is "6 years": "June 1994" $\rightarrow$ "June 2000"; "2008" $\rightarrow$ "2014"). For each relation, we generate these random names with LLMs by providing several examples of existing entity names. To avoid confusing the model, we only use obscure names that do not exist in our dataset.
$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e)$ (6)
$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c))$
$a^{\prime\prime}=f_{T}(f_{E}(a))$
where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity name changing and time changing, respectively.
4 Experiments
We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation? (2) Can our two-step framework lead to better temporal reasoning performance? (3) Can LLMs benefit from our data on other reasoning tasks?
4.1 Experimental Setup
In addition to TempGraph QA, we consider the existing dataset, TimeQA  (Chen et al., 2021), which is based on the texts from Wikipedia. It contains two difficulty levels: The easy-level split tends to be information extraction while the hard-level split involves temporal reasoning, i.e., understanding the relation between different temporal expressions. We evaluate our framework on both datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider the accuracy based on perplexity for LLM evaluations. Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format. This property actually hurts their performance under the commonly used metrics such as EM and F1. To solve this problem, we select the answers from a candidate set based on their perplexity from the LLM. For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them. For TempGraph QA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
We compare our framework with leading LLMs including Llama2, GPT-3.5 and GPT-4. Llama2 Touvron et al. (2023b) is one of the most accepted open-source LLMs, while GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023) are two powerful closed-source models. We run Llama2 series on local machines (4-bit quantization is adopted for 70B version due to computational resource limitation), and use APIs provided by OpenAI for GPT series. We evaluate their few-shot ICL performance on the test set. To avoid LLMs generating free text which affects the evaluation with these metrics  (Yang et al., 2023a), we provide candidate answers for all questions under the ICL setting. The model versions and prompt templates are shown in Appendix B and C, respectively.
4.2 Implementation Details
We use Llama2-13B as the baseline due to computation resource limit. In experiments, we proved that our framework has comparable or even better temporal reasoning performance than one of the most powerful LLMs, GPT-4. We inject two adapters with selectors into the Llama2 model for the text-to-tempgraph translation and symbolic temporal reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform symbolic temporal reasoning on it, i.e., the adapters are used in sequence. In addition, we use GPT-3.5 for the story creation & verification in TempGraph QA, entity and relation extraction, temporal info identification, graph construction & verification on other datasets, and symbolic CoTs bootstrapping. These examples used in ICL are created by GPT-4 with manual verification. We choose these models due to the high quality of their generation. All the prompt templates are shown in Appendix C.
4.3 Main Results
Can our strategies guide LLMs to perform more reliable symbolic reasoning?   Figure 2 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TempGraph QA. To obtain a fair comparison, all the reasoning happens on the estimated graphs obtained from our text-to-graph translation module. It can be seen that LLMs learn better temporal reasoning with SFT than ICL. By providing CoTs with symbolic reasoning, the model performance gets further enhanced. There is an additional performance boost when using our data augmentation strategies. Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies. Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error). We measure the level of agreement among the three evaluators, and show the error rates from different types in Table 4. It can be seen that LLMs with ICL easily use either wrong information or erroneous external knowledge in reasoning. This issue can be alleviated with SFT, and we shows that using symbolic reasoning is better than free-text CoTs. Our data augmentation strategies further reduce all types of errors, and provides a promising direction for LLMs on reasoning tasks.
Figure 2: Performance comparison between different strategies of CoT generation on TempGraph QA.
Strategy ER-T1 ER-T2 ER-T3 ER-T4
ICL 0.13 0.13 0.31 0.10
SFT 0.04 0.17 0.13 0.10
SFT + bs 0.06 0.13 0.03 0.08
SFT + bs + aug 0.03 0.05 0.04 0.05
Table 4: Human evaluation on the CoTs generated by different strategies. T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error.
Model TempGraph QA TimeQA
Easy-mode Hard-mode
EM F1 Acc EM F1 Acc EM F1 Acc
GPT-3.5 (ICL-SP) 0.559 0.699 - 0.640 0.668 - 0.512 0.506 -
GPT-3.5 (ICL-CoT) 0.698 0.788 - 0.565 0.603 - 0.436 0.464 -
GPT-4 (ICL-SP) 0.704 0.818 - 0.716 0.742 - 0.571 0.546 -
$*{}^{*}$ (ICL-CoT) 0.798 0.833 - 0.662 0.693 - 0.618 0.636 -
Llama2-7B (ICL-SP) 0.402 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354
Llama2-7B (ICL-CoT) 0.548 0.686 0.578 0.367 0.425 0.393 0.302 0.354 0.360
Llama2-13B (ICL-SP) 0.423 0.609 0.526 0.439 0.493 0.450 0.427 0.481 0.437
Llama2-13B (ICL-CoT) 0.626 0.762 0.668 0.518 0.572 0.535 0.434 0.490 0.480
Llama2-70B (ICL-SP) 0.578 0.736 0.665 0.583 0.627 0.631 0.493 0.537 0.551
Llama2-70B (ICL-CoT) 0.760 0.838 0.796 0.552 0.612 0.623 0.447 0.501 0.512
Llama2-13B (SFT-SP) 0.550 0.720 0.652 0.412 0.449 0.455 0.362 0.404 0.412
Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379
Llama2-13B (TempGraph) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649
Table 5: Main results using different models and strategies. We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc). Results with * are evaluated on 1000 random test samples.
Figure 3: Performance comparison between different strategies on TimeQA. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies.
Can our two-step framework lead to better temporal reasoning performance?   Table 5 shows the comparison between different models on both TempGraph QA and TimeQA. First, we observe that among all the LLMs with ICL, GPT-4 has the best temporal reasoning capability as expected. For Llama2 series, larger model has better performance due to advanced context understanding and improved generalization. We also find that CoTs not always bring better temporal reasoning. In both easy-mode and hard mode of TimeQA, CoTs hurt larger LMs such as GPT-3.5, GPT-4 and Llama2-70B due to unreliable intermediate results and hallucinations. More importantly, our model, which is based on Llama2-13B, shows a comparable or even better temporal reasoning performance than GPT-4 on both datasets. To prove the effectiveness of our framework, we also provide the results of some alternative strategies (ICL, SFT-SP, SFT-CoT), where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively. Compared with our framework, there are no text-to-graph translation and symbolic reasoning involved in SFT-SP and SFT-CoT. It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets. Compared with original text, we actually provide an easier path from temporal reasoning questions to their answers.
Can LLMs benefit from our data on other reasoning tasks? Figure 3 shows the performance comparison between different strategies on TimeQA (ICL-SP, ICL-CoT, pre-training on TempGraph QA, SFT on TimeQA). It can be seen that the performance of in-context learning (either with standard I/O prompt or CoTs) is much lower than pre-training and SFT. In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning. More importantly, we observe that per-training on synthetic data improves the model performance compared with ICL, which proves the rationality of design and construction of our dataset. Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions. Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the reasoning capability of LLMs.
4.4 Ablation Study
We ablate different modules to see their contributions to the model performance. Table 6 shows the performance comparison between different configurations where TG, EK, SR denote temporal graph, external knowledge and symbolic reasoning, respectively. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. We use supervised fine-tuning on CoTs (SFT-CoT) as baseline. We improve the model performance 2.9% in EM, 3.2% in F1 and 0.6% in Acc by adding TG. Then we show that bootstrapping CoTs with symbolic reasoning brings further enhancement of 4.8% in EM, 4.3% in F1 and 2.5% in Acc. These is a significant boost in EM (5.9%), F1 (4.1%) and Acc (5.9%), when using our data augmentation strategies. Finally, the introduction of external knowledge exhibits meaningful results: it increases EM by 3%/1.5%, F1 by 0.6%/1.2%, and Acc by 4%/2.4% without/with data augmentation. From the ablation study, we give some insights for LLMs: (1) LLMs can benefit from (temporal) graph which is intuitive, concise and structured. (2) Given a reliable graph, symbolic reasoning brings better performance in reasoning than free-text CoTs. (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning. (4) The introduction of external knowledge such as mathematics and commonsense can augment the generation of LLMs.
Config EM F1 Acc
Baseline (SFT-CoT) 0.646 0.722 0.705
TG 0.675 0.754 0.711
TG + SR(bs) 0.723 0.797 0.736
TG + SR(bs + aug) 0.782 0.838 0.795
TG + EK + SR(bs) 0.753 0.803 0.776
TG + EK + SR(bs + aug) 0.797 0.850 0.819
Table 6: Ablation Study Results on TempGraph QA. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
5 Related Work
Text-based Temporal Reasoning   Recently, text-based temporal reasoning has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023). The vision here is to help LMs understand temporal concepts, which are fundamental elements in human knowledge. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust temporal reasoning.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge or extra temporal constraints. Although these methods make some progress, the underlying structure and logic of temporal reasoning are either ignored or not explicitly involved. Since LMs perform pre-training on human knowledge, the manual perspective of modelling might effectively improve their performance on this task.
Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and their capability of reasoning (Valmeekam et al., 2022; Wei et al., 2022; Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modelling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing the final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that requires reasoning. To the best of our knowledge, the exploration of the full capabilities of LLMs on temporal reasoning is still missing. Compared with other reasoning types, it is a form of integrated task that requires the knowledge of mathematics, logic and commonsense.
Neuro-symbolic Reasoning   Recent research shows advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Susskind et al., 2021; Hamilton et al., 2022). Knowledge graph is an important representation as the foundation of symbolic reasoning. Existing works on symbolic reasoning over KGs include  (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021). On top of that,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between text-based and temporal KG-based temporal symbolic reasoning. This connection brings the potential to extend these KG-based methods on text-based tasks. Some examples include  (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023). Different from these methods, which limit their application to certain datasets, our framework has better generalization due to the text-to-graph translation part.
6 Conclusion
TempGraph-LLM, a new paradigm for language models, has been proposed to improve their performance on temporal reasoning. Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning process. Experiments indicate that TempGraph-LLM achieves better performance than existing methods. An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning. Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks with good problem formalization and methodology extension.
Appendix A Dataset Statistics
In this section, we show statistics for our TempGraph QA dataset. Figure 4 shows the number of facts in training and test set, which determine the complexity of task to some extent. Figure 5 shows the types of questions in training and test set. Different types of questions are exemplified in Table 1. Among all the types, question 1 has the largest portion since it has five variants. Similarly, question 7 with two variants also has a larger portion. To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores.
Figure 4: Number of facts in TempGraph QA.
Figure 5: Types of questions in TempGraph QA.
Appendix B Model Versions
All the model versions used in our experiments are listed below:
  • Llama2-7B (meta-llama/Llama-2-7b-hf)
  • Llama2-13B (meta-llama/Llama-2-13b-hf)
  • Llama2-70B (meta-llama/Llama-2-70b-hf)
  • GPT-3.5 (gpt-3.5-turbo)
  • GPT-4 (gpt-4-1106-preview)
Appendix C Example Prompts
In this section, we show example prompts used in our framework. To be specific, Table 7 shows an example of the graph-based story generation in TempGraph QA; Table 8 presents an example of the automatic story-tempgraph alignment verification in TempGraph QA; Table 9 provides an example of the temporal info identification in TimeQA; Table 10 illustrates an example of the graph construction in TimeQA; Table 11 gives an example of the automatic tempgraph-QA alignment verification in TimeQA; Table 12 depicts an illustrative case of the CoTs bootstrapping in TempGraph QA.
(Molly Adams was born in Seattle) starts at 1896
(Liam Thomas Dawson was born in Seattle) starts at 1896
(Molly Adams was married to Liam Thomas Dawson) starts at 1920
(Liam Thomas Dawson was married to Molly Adams) starts at 1920
(Liam Thomas Dawson won prize James Parker Prize for Advanced Biomedical
Studies) starts at 1946
(Liam Thomas Dawson won prize Champion Award in Biology or Science) starts at 1947
(Liam Thomas Dawson won prize Oakley Smith Prize) starts at 1948
(Liam Thomas Dawson won prize Member of the National Academy) starts at 1950
(Molly Adams was married to Liam Thomas Dawson) ends at 1957
(Liam Thomas Dawson was married to Molly Adams) ends at 1957
(Liam Thomas Dawson died in Kingsville, Texas) starts at 1984
Based on the above facts, generate a short story for me.
Table 7: Example of the graph-based story generation in TempGraph QA.
Once upon a time, in the vibrant city of Seattle, two remarkable individuals, Molly
Adams and Liam Thomas Dawson, were born in the year 1896. Little did they know that
their destinies were entwined from the very beginning.
Growing up in the same neighborhood, Molly and Liam developed a deep friendship
that blossomed into something more as they entered adulthood. In the year 1920, their
love story officially began as they exchanged vows and embarked on a journey of
companionship that would last for decades.
$\cdots$
When did the event (Molly Adams was married to Liam Thomas Dawson) end?
Table 8: Example of the automatic story-tempgraph alignment verification in TempGraph QA.
Knox Cunningham
Sir Samuel Knox Cunningham, 1st Baronet, QC (3 April 1909 – 29 July 1976) was a
Northern Irish barrister, businessman and politician. As an Ulster Unionist politician at
a time when the Unionists were part of the Conservative Party, he was also a significant
figure in United Kingdom politics as Parliamentary Private Secretary to Harold
Macmillan. His nephew was Sir Josias Cunningham.
$\cdots$
Extract all the time expressions such as ’June 1994’, ’1973’, ’late 1980s’.
Table 9: Example of the temporal info identification in TimeQA.
Knox Cunningham
Sir Samuel Knox Cunningham, 1st Baronet, QC (3 April 1909 – 29 July 1976) was a
Northern Irish barrister, businessman and politician. As an Ulster Unionist politician at
a time when the Unionists were part of the Conservative Party, he was also a significant
figure in United Kingdom politics as Parliamentary Private Secretary to Harold
Macmillan. His nephew was Sir Josias Cunningham.
$\cdots$
Construct a timeline for Knox Cunningham’s position. You should only consider these
time points (3 April 1909, 1930s, 1939, 1942, 1943, 1945, 1947, 1949, 1954, 29 July 1976).
Table 10: Example of the graph construction in TimeQA.
3 April 1909: Knox Cunningham was born.
1930s: He studied law and began his legal career.
1935: On 2 July 1935, he married Dorothy Enid Riley JP.
1939: Cunningham was called to the Bar by the Middle Temple.
1942: He was called to the Bar in Northern Ireland.
1943: Cunningham fought the Belfast West by-election.
1945: He contested the same seat in the general election.
1947: Knox Cunningham became involved with the World Alliance of YMCAs.
1949: He became Chairman of the National Council of the YMCA.
1954: Cunningham was elected to Orpington Urban District Council.
29 July 1976: Knox Cunningham passed away.
Which position did Knox Cunningham hold from May 1955 to Apr 1956?
Table 11: Example of the automatic tempgraph-QA alignment verification in TimeQA.
Example:
Timeline:
(Liam Mitchell was born in Harrison) starts at 1885
(Maddox Reynolds was born in Glenville) starts at 1893
(Liam Mitchell was married to Emma Scott) starts at 1916
(Emma Scott was married to Liam Mitchell) starts at 1916
(Liam Mitchell was married to Emma Scott) ends at 1918
(Emma Scott was married to Liam Mitchell) ends at 1918
(Maddox Reynolds was married to Liam Mitchell) starts at 1922
(Liam Mitchell was married to Maddox Reynolds) starts at 1922
(Maddox Reynolds was married to Liam Mitchell) ends at 1928
(Liam Mitchell was married to Maddox Reynolds) ends at 1928
(Liam Mitchell died in Boston) starts at 1941
(Emma Scott died in Oceanview) starts at 1984
Question: True or false: event (Liam Mitchell was married to Maddox Reynolds) was
longer in duration than event (Emma Scott was married to Liam Mitchell)?
Useful information:
1885 before 1893 before 1916 before 1918 before 1922 before 1928 before 1941
1918 - 1916 = 2
1928 - 1893 = 35
1928 - 1922 = 6
1941 - 1918 = 23
2 < 6 < 23 < 35
Answer: Let’s think step by step.
To determine whether the statement is true or false, we need to compare the durations
of (Liam Mitchell was married to Maddox Reynolds) and (Emma Scott was married to
Liam Mitchell) . The duration for each event can be calculated as follows:
(Liam Mitchell was married to Maddox Reynolds) :
(Liam Mitchell was married to Maddox Reynolds) starts at 1922
(Liam Mitchell was married to Maddox Reynolds) ends at 1928
1928 - 1922 = 6
(Emma Scott was married to Liam Mitchell) :
(Emma Scott was married to Liam Mitchell) starts at 1916
(Emma Scott was married to Liam Mitchell) ends at 1918
1918 - 1916 = 2
Now, we compare the duration of the two events: 6 is greater than 2 , thus, the answer is True .
Test:
Timeline:
(Chris Evans was born in Bristol, Connecticut) starts at 1948
(Oliver Jensen was born in Bristol, Connecticut) starts at 1948
(Chris Evans was married to Oliver Jensen) starts at 1970
(Oliver Jensen was married to Chris Evans) starts at 1970
(Chris Evans created Pearl Network) starts at 2005
(Chris Evans won prize Victory Achievement Award) starts at 2007
(Chris Evans was married to Oliver Jensen) ends at 2010
(Oliver Jensen was married to Chris Evans) ends at 2010
(Chris Evans created Pearl Network) ends at 2013
Question: True or false: event (Chris Evans owned Pearl Network) was longer in
duration than event (Oliver Jensen was married to Chris Evans)?
Useful information:
1948 before 1970 before 2005 before 2007 before 2010
1970 - 1948 = 22
2007 - 1948 = 59
2007 - 2005 = 2
2010 - 1970 = 40
2 < 22 < 40 < 59
Answer: Let’s think step by step.
Table 12: Example of the CoTs bootstrapping in TempGraph QA.

### v2

@@
 Abstract
-Large language models (LLMs) learn temporal concepts from the co-occurrence of related tokens in a sequence.
-Compared with conventional text generation, temporal reasoning, which reaches a conclusion based on mathematical, logical and commonsense knowledge, is more challenging.
-In this paper, we propose TempGraph-LLM, a new paradigm towards text-based temporal reasoning.
-To be specific, we first teach LLMs to translate the context into a temporal graph.
-A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for pre-training on this task.
-We prove in experiments that LLMs benefit from the pre-training on other tasks.
-On top of that, we guide LLMs to perform symbolic reasoning with the strategies of Chain of Thoughts (CoTs) bootstrapping and special data augmentation.
-We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free text.
+While large language models (LLMs) have demonstrated remarkable reasoning capabilities, they are not without their flaws and inaccuracies.
+Recent studies have introduced various methods to mitigate these limitations.
+Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate contextual details.
+In this paper, we propose TG-LLM, a new framework towards language-based TR.
+To be specific, we first teach LLM to translate the context into a temporal graph (TG).
+A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for fine-tuning on this graph translation task.
+We confirm in experiments that the capability of TG extraction learned on our dataset can be transferred to other TR tasks and benchmarks.
+On top of that, we guide LLM to perform symbolic reasoning over the TG via Chain of Thoughts (CoTs) bootstrapping and special data augmentation strategies.
+We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free-form text.
 1 Introduction
-As one of the fundamental abilities, temporal reasoning plays an important role in human perception.
+As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception.
 It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations.
-Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023).
-However, they still can not perform temporal reasoning very well as we expected  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023).
-Compared with other singular types, temporal reasoning is a form of integrated task built upon foundations like mathematics, logical and commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
-Figure 1: Our TempGraph-LLM framework has two steps: text-to-tempgraph translation and neuro-symbolic reasoning.
-Recent work mainly adopts the general methods to investigate and improve the temporal reasoning capability of LLMs.
-For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different temporal reasoning tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).
+Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023).
+However, they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023).
+TR is more challenging since it requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
+Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and neuro-symbolic reasoning.
+Recent works mainly adopt the general reasoning methods to investigate and improve the TR capability of LLMs.
+For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different TR tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).
 (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT).
 Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL).
-In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include span extraction pre-training, instruction tuning and reinforcement learning.
-Despite the effectiveness of such methods, they ignore the intrinsic nature of temporal reasoning, and cannot guarantee the correctness of intermediate steps  Saparov and He (2023); Yang et al.
-(2023b).
-Human performs temporal reasoning on a timeline of events which are aligned with the entities and relations.
+In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
+Despite the effectiveness of such methods, they either ignore or not explicitly involve the intrinsic nature of TR.
+Human performs complex TR on a timeline of events which are aligned with the entities and relations.
 These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information.
-Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline from the text, and then to perform deliberate reasoning on it.
-However, the high-quality extraction of aligned timeline is a challenging task itself due to lack of ground truth.
+Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline of events from the text, and then perform deliberate reasoning on it.
+However, the high-quality extraction of the aligned timeline is a challenging task itself due to the lack of ground truth.
 To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision.
-We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks.
-Given the reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on that.
-(1) How to introduce the necessary arithmetic and commonsense knowledge involved in temporal reasoning?
-(2) How to guide LLMs to perform rigorous reasoning?
-Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap of the timestamps, the relative order of the time gaps).
-As for (2), there are always two roadmaps: (i) translating natural language into logical statements and using symbolic inference engine for reasoning  Pan et al.
+We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks and benchmarks.
+Given a reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on it: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR?
+(2) How can one guide LLMs to perform rigorous reasoning?
+Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps).
+As for (2), there are two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al.
 (2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023).
@@
 (2023c) and the limited expressive power of formal logic.
-For (ii), there is no guarantee for the generated text especially when training data are insufficient  (Yang et al., 2023b).
-To solve these problems, we propose a symbolic integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning.
-Furthermore, we design two types of data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
+For (ii), there is no guarantee for the correctness of the generated text especially when training data are insufficient  (Yang et al., 2023b).
+To solve these problems, we propose a symbolic-integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning.
+Furthermore, we design special data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
 To be specific, our contributions are summarized as follows:
-  • We propose a new paradigm, TempGraph-LLM, for text-based temporal reasoning.
-In this framework, we first translate the context into a temporal graph, and then perform symbolic reasoning on it.
-Extensive experiments prove that it brings better performance than convetional methods.
-  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLMs to utilize symbolic reasoning, which brings better performance than free text.
-  • We present a pipeline to create a synthetic dataet for question answering that requires temporal reasoning.
-It is fully controllable and requires minimal supervision for context-tempgraph alignment.
-We show in experiments that pre-training on our dataset benefits LLMs on other tasks.
-2 Dataset Construction
-We propose a pipeline for the construction of our synthetic temporal QA dataset called TempGraph QA.
-Each sample, as shown in Table 1, is in the form of (story, tempgraph, question, answer).
-The tempgraph is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018).
-The number of events in each tempgraph is restricted into a certain range.
-By using some rule-based templates, we obtain reliable question and answer pairs from the graph.
-Finally, each story is translated from the tempgraph by GPT-3.5.
-Below we show the details for all steps.
-TempGraph (sub; rel; obj; start/end; time):
+  • We propose a new paradigm, TG-LLM, for language-based TR.
+In this framework, we first translate the context into a temporal knowledge graph, and then perform symbolic reasoning on it.
+Extensive experiments prove that our novel approach results in superior performance compared to the conventional approach.
+  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLM to utilize symbolic reasoning, which brings better performance than free-form text.
+  • We present a pipeline to create a synthetic dataset for question answering that requires TR.
+It is fully controllable and requires minimal supervision for text-temporal graph alignment.
+We show in experiments that fine-tuning on our dataset benefits LLMs on other TR tasks and benchmarks.
+Temporal Graph (sub; rel; obj; start/end; time):
 […]
@@
 and an adventurous soul. $\cdots$
-Graph-based QA (from Rule-based Python Script):
+Graph-based QAs (from rule-based Python script):
 […]
@@
 Sophia Parker)?
-A: (John Thompson owned Pearl Network).
-Q: True or false: event (John Thompson owned Pearl
+A1: (John Thompson owned Pearl Network).
+Q2: True or false: event (John Thompson owned Pearl
 Network) was longer in duration than event (Sophia
 Parker was married to John Thompson)?
-A: True.
+A2: True.
 $\cdots$
-Table 1: Each sample from TempGraph QA is in the form of (story, tempgraph, question, answer).
+Table 1: Each sample from TGQA dataset is in the form of (story, temporal graph, questions, answers).
+2 Dataset Construction
 Reasoning Type Question Answer
-Sequencing Which event occurred first, {event A} or {event B}? {event A}/{event B}
-Given the following five events: {event A}, {event B}, {event C}, {event D}, {event A}/{event B}/
-{event E}, which event is the first/second/third/fourth/fifth one in chronolog- {event C}/{event D}/
-ical order? {event E}
-Duration How long did the {event A} last? {duration of event A}
-Temporal How much time passed between the start of {event A} and the {gap between event A
-Relation start of {event B}? and B’s start time}
-What happened right before/after {event A} started? {event B} right before/
-after {event A}
-Facts When did the {event A} occur? {event A’s start time}
+Sequencing Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_A>/<Event_B>/
+<Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in <Event_C>/<Event_D>/
+the chronological order? $\cdots$
+Duration How long did the event <Event_A> last? <Duration_of_Event_A>
+Temporal How much time passed between the start of <Event_A> and the start of <Gap_between_Event_A
+Relation <Event_B>? _and_B_start_time>
+What happened right before/after <Event_A> started? <Event_B> right before/
+after <Event_A>
+Facts When did the <Event_A> occur? <Event_A_start_time>
 Extraction
-Simultaneity True or false: {event A} and {event B} happened at the same year? True / False
-True or false: {event A} was still happening when {event B} started? True / False
-Comparative True or false: {event A} was longer in duration than {event B}? True / False
+Simultaneity True or false: <Event_A> and <Event_B> happened at the same year? True / False
+True or false: <Event_A> was still happening when <Event_B> started? True / False
+Comparative True or false: <Event_A> was longer in duration than <Event_B>? True / False
 Analysis
-Table 2: Reasoning types with corresponding questions and answers in TempGraph QA.
-Step 1: Graph Splitting & Masking   Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them all at once.
+Table 2: Reasoning types with corresponding questions and answers in TGQA.
+We propose a pipeline for the construction of our synthetic dataset called TGQA.
+Each sample, as shown in Table 1, is in the form of (story, temporal graph, questions, answers).
+The temporal graph (TG) is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), with a restriction on the number of events.
+Each story is translated from the TG by GPT-3.5.
+Finally, by using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph.
+We obtain 500 samples for training/validation and another 100 for test, with about 30 QA pairs in a single sample.
+Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of temporal questions as shown in Table 2.
+Below we provide the details of each step in the dataset construction pipeline.
+Step 1: Graph Splitting & Masking  Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them into text all at once.
 Given a specific entity, we extract its third-order neighbors and all the events happening between them.
-To avoid the discrepancy caused by the number of events, we set a certain range and discard the non-compliant subgraphs.
-Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training and test samples.
-We finally obtain 500 tempgraphs for training and another 100 for test.
-In addition, we consider the noticeable leakage problem of LLMs, where the prior knowledge of test data has been implicitly obtained from the extensive pre-training.
-Thus, a masking strategy, i.e., changing entities names into random ones of the same type, is adopted.
-We generate a global mapping for entity names with GPT-3.5.
+Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training/validation and test samples.
+In addition, we consider the noticeable corpus leakage problem of LLMs, where prior knowledge of test data has been implicitly obtained from the extensive pre-training process.
+Thus, a masking strategy, i.e., changing entity names into random ones of the same type, is adopted.
+We generate a global mapping for entity names with GPT-3.5 (see Table 6).
 To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
-Step 2: Graph-based Open QA Creation   Based on the given tempgraph, we generate a story and multiple QAs.
+Step 2: Graph-based Open QA Creation  Based on the given TG, we generate a story and multiple QAs.
 We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts.
 It is observed that GPT-3.5 tends to ignore the end time of some events in the created story.
-To solve this problem, we split the starting and ending of a persistent event into two transient ones.
+To solve this problem, we split the starting and ending of a persistent event into two transient ones (see Table 1).
 To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction.
-Given these categories, we design nine different question types (shown in Table 2), and write a rule-based Python script to determine the answers from the tempgraph.
-Step 3: Quality Control   Although LLMs are good at story generation, there is a risk of misalignment between the given tempgraph and the generation text.
-To avoid the interference caused by misalignment, we propose an automatic way for verification.
+Given these categories, we design multiple diverse question types (see Table 2), and use a rule-based Python script to generate the corresponding Qs and As.
+Step 3: Quality Control  Although LLMs are good at story generation, there is a risk of misalignment between the given TG and the generated text.
+To avoid the interference caused by misalignment, we propose a semi-automatic way for verification.
 Given a generated story, we ask LLMs the time of each event in the graph.
-If LLMs cannot give the correct answer, we remove the possibly missing events in the graph.
-Since this strategy tends to simplify the reasoning, we manually verify the existence of the possibly missing events for both training and test sets.
-Note that only the quality of story-tempgraph alignment needs control, since all the QAs are generated from the tempgraph with specific rules.
-We show dataset statistcis in Appendix A, and all the prompts involved in Appendix C.
-3 TempGraph-LLM
-Motivated by human perception, we propose a new paradigm called TempGraph-LLM. We first translate the text into a temporal graph, and then guide LLMs to perform symbolic reasoning on it.
-3.1 Text-to-TempGraph Translation
-TempGraph (TG) serves as the foundation of the following symbolic reasoning process.
-Although LLMs can extract them from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs.
+If LLMs cannot give the correct answer, we consider the event as possibly missing in the story.
+Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors and only manually inspect the set of unanswered questions by LLM.
+Note that only the quality of story-TG alignment needs manual inspection, since all the QAs are generated from the TG with specific rules.
+We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
+3 TG-LLM
+Motivated by human perception, we propose a new paradigm called TG-LLM. We first translate the text into a temporal graph, and then guide the LLM to perform symbolic reasoning on it.
+3.1 Text-to-TG Translation
+TG serves as the foundation of the following symbolic reasoning process.
+Although LLMs can extract fact/events from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs, i.e, LLMs making mistakes or using different entity/relation names or time expressions.
 Thus, we provide a deliberate pipeline for high-quality TG translation.
-3.1.1 Pre-training
-In the TempGraph QA, each story is corresponding to a verified TG.
-Similarly, for other data, where the ground truth of TG is avaiable, we can pre-train LLMs on them.
-Since language models (LMs) can only deal with sequence of tokens, we sort all the events in the graph according to their time.
-In addition, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events.
+Ground-truth TG Generation   For datasets such as TGQA, each story is corresponding to a verified TG.
+Thus, we can directly fine-tune LLMs on this task with the ground truth.
+However, the biggest challenge for most real-world applications is the lack of ground truth.
+Instead of using the naive estimation from LLMs, we provide a pipeline for high-quality generation.
+(1) Entity & Relation Extraction: To align the entity/relation names in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context.
+To obtain the entity/relation names in pre-defined QAs, we consider two ways: parsing with rules or using LLMs.
+(2) Temporal Info Identification: To identify all the time expressions, LLMs are first used to extract them from the context, and then a rule-based Python script is adopted for time normalization and invalid output filtering.
+For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, which are introduced as external knowledge in temporal reasoning (see Table 7).
+(3) Graph Construction & Verification: We generate the graph using LLMs with ICL.
+To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (see Table 15).
+For those without explicit expressions, we ask LLMs to provide the temporal order of different events.
+We further semi-automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph.
+For those graphs where LLMs give the wrong answers, we manually check and correct them.
+Figure 2: In chain of thoughts (CoTs) bootstrapping, we accept CoTs leading to correct answers and sample them according to their contrastive learning scores.
+Supervised Fine-tuning  We found out in our experiments that providing a chronological list of events helps LLMs provide better temporal reasoning.
+In addition, as shown in Table 1, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events.
 With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times.
-We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023b) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021).
+We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021).
 The input and output of the model are the story and aligned timeline, respectively.
-In experiments, we observe benefits from the pre-training on our data for other tasks.
-3.1.2 Fine-tuning
-Generally, the biggest challenge for fine-tuning on real-world datasets is lack of ground truth. Instead of directly using the estimation from LLMs, we provide a pipeline for better alignment.
-Entity & Relation Extraction   To align the entities and relations in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context.
-To obtain the entities and relations in pre-defined QAs, we consider two ways: parsing the QAs with rules or using LLMs.
-The former is more efficient and reliable when the QAs are generated by templates  (Chen et al., 2021; Tan et al., 2023a), while the latter should be used when the QAs are in free text  (Zhou et al., 2019; Ning et al., 2020).
-Temporal Info Identification   Compared with conventional reasoning tasks, temporal information is one of the most important and unique components here.
-To identify all the time expressions, LLMs are first used to extract them in the context, and then a rule-based Python script  (Yang et al., 2023a) is adopted for time-normalization and invalid output filtering.
-For example, we convert "Dec.
-5th, 2004" into "2004-12-05", "July 1989" into "1989-07-XX", and "1973" into "1973-XX-XX", respectively.
-For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, e.g., "1974 is before 1983", "the time gap between May 2004 and Oct 2008 is 4 years and 5 months", which are introduced in temporal reasoning as external knowledge.
-Graph Construction & Verification   We generate the graph using LLMs with in-context learning (ICL).
-To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output.
-The TG should only contain the extracted elements to guarantee the alignment with pre-defined QAs.
-We classify the stories according to whether they contain explicit expressions such as "July 1989", "2003–2005", "the late 1900s".
-For those without explicit expressions, we ask LLMs to provide the temporal order of different events.
-We further automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph.
-For those graphs where LLMs give the wrong answers, we manually check and correct them.
+In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
 3.2 Neuro-symbolic Temporal Reasoning
-Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by Chain of Thoughts (CoTs) bootstrapping and data augmentation.
+Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by CoTs bootstrapping and data augmentation.
 3.2.1 Bootstrapping Chain of Thoughts
@@
 Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples.
-We bootstrap the generations and only accept the ones with correct final answers.
-To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{m}\}$ with LLMs.
-For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask LLMs to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$.
+To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{M}\}$.
+For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask GPT-3.5/GPT-4 to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$.
 These CoTs will be manually verified and used as few-shot ICL examples.
-When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c^{k}_{j^{\prime}},\hat{a}^{k}_{j^{\prime}})$ from LLMs, and only accept the one leading to correct answers, i.e., $\hat{a}^{k}_{j^{\prime}}=a^{*}_{j^{\prime}}$.
-Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method is more flexible and robust.
-In addition, inspired by  (Wang et al., 2023), we compare different type of CoTs using contrastive learning.
-The core idea is to choose CoTs that make the correct answer more plausible than other candidates.
-$c^{*}=\operatorname{argmax}\log P\left(a^{*}\mid g,e,q,c\right)+\gamma G\left(c\right)$ (1)
-$G\left(c\right)=\log\frac{P\left(a^{*}\mid g,e,q,c\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c\right)}$ (2)
-$P\left(a\mid g,e,q,c\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c,t_{<l}\right)$ (3)
-where $g,e,q,c$ denote the TG, external knowledge, question, and CoTs, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
+When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})$ from GPT-3.5, and only accept the one leading to correct answers, i.e., $\hat{a}_{j^{\prime},k}=a^{*}_{j^{\prime}}$.
+In addition, inspired by  (Wang et al., 2023), we compare different CoTs using contrastive learning given in Eq.
+2.
+The core idea is to choose CoTs that make the correct answer more plausible than other candidate answers.
+For the generated CoTs, we use weighted sampling according to the score-based probability given in Eq.
+1.
+$P_{sample}(c_{k})=\operatorname{softmax}(score(c_{k}))$ (1)
+$score(c_{k})=\log P\left(a^{*}\mid g,e,q,c_{k}\right)+\gamma G\left(c_{k}\right)$ (2)
+$G\left(c_{k}\right)=\log\frac{P\left(a^{*}\mid g,e,q,c_{k}\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c_{k}\right)}$ (3)
+$P\left(a\mid g,e,q,c_{k}\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c_{k},t_{<l}\right)$ (4)
+where $P_{sample}(\cdot)$ is the sampling probability for CoTs selection, $score(\cdot)$ is the contrastive learning score for each CoT, $g,e,q,c$ denote the TG, external knowledge, question, and CoT, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
+Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method has more diversity and does not require prior knowledge for the reasoning process.
 3.2.2 Data Augmentation
 Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023).
-To address the issue, we propose two special strategies for data augmentation.
-Learning with Disturbances   Essentially, our framework involves two steps: text-to-tempgraph translation and symbolic temporal reasoning.
-At the test stage, reasoning happens on the noisy graphs, which can impact the overall performance.
-Specifically, for the reasoning module, we train with ground truth/verified graphs while testing on the estimations.
+To address the issue, we propose special strategies for data augmentation.
+Essentially, our framework involves two steps: text-to-TG translation and symbolic temporal reasoning.
+At the test stage, reasoning happens on the noisy estimated graphs, which impacts the overall performance.
+Specifically, for the reasoning module, we train with ground-truth/verified graphs while in the test time, the estimated graph is used for symbolic reasoning.
 To improve the robustness of reasoning, we introduce some disturbances on the graphs during training.
 Note that the type of disturbances should be carefully designed in order to avoid confusing the model.
-We investigate two types of disturbances: remove irrelevant edges or replace edges by changing relation names, e.g., "owned" $\rightarrow$ "run", "was affiliated to" $\rightarrow$ "belonged to", "played for" $\rightarrow$ "joined".
-An edge is considered as irrelevant if it is not involved in both QA and CoTs.
+We first investigate two types of disturbances: (i) remove irrelevant edges and (ii) replace edges by changing relation names, given in Eq.
+5 and Eq.
+6, respectively.
+An edge is considered irrelevant if it is not involved in both QA and CoT.
 $F_{irr}\in\{F:P(c|g_{\setminus\{F\}},e,q)\approx P_{0}\}_{F\in g}$ (5)
 $F^{\prime}\in\{f_{R}(F):P(c|f_{R}(g),e,q)\approx P_{0}\}_{F\in g}$ (6)
-TempGraph (Adding Disturbances):
+Figure 3: We have several strategies for data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
+where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation names $f_{R}(\cdot)$.
+Furthermore, to make sure LLM learns the underlying principle of symbolic temporal reasoning instead of just memorizing the training data, we introduce another two types of disturbances: (iii) globally map all the entity names in training data to some random names of the same type , and (iv) change the times based on a global offset, given in Eq.
+7.
+For each relation, we generate these random names with GPT-3.5 by providing several examples of existing names.
+$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e),$ (7)
+$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c)),$
+$a^{\prime\prime}=f_{T}(f_{E}(a))$
+where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity names and time changing, respectively.
+4 Experiments
+We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation?
+(2) Can our two-step framework lead to better temporal reasoning performance?
+(3) Does these learned capabilities of temporal reasoning generalize to other tasks?
+4.1 Experimental Setup
+In addition to TGQA, we consider the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries.
+Compared with these datasets, TGQA has a better quality due to the controlled generation, and more diverse question categories (see examples from other datasets in Appendix B).
+We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
+Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations, i.e., selecting from a candidate set the answer with the lowest perplexity from LLMs as their prediction.
+(See rationales for evaluation metrics selection in Appendix B)
+We compare our framework [^1] Code and data available at https://github.com/xiongsi heng/TG-LLM.
+with leading LLMs including Llama2 Touvron et al.
+(2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023).
+We run the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and use APIs provided by OpenAI for GPT models.
+We evaluate their few-shot ICL performance on the test set.
+To avoid LLMs generating free text which affects the evaluation, we provide candidate answers for all questions.
+The model versions and prompt templates are provided in Appendix B and C, respectively.
+We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison.
+Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
+4.2 Implementation Details
+We use Llama2-13B as the baseline due to limited computational resources.
+We inject two adapters with selectors into the model for the text-to-TG translation and symbolic temporal reasoning.
+The adapters are trained in parallel.
+For inference, we first translate the original story into a temporal graph, and then perform symbolic reasoning on it, i.e., the adapters are used in sequence.
+In addition, we use GPT-3.5/GPT-4 for all the generation of story, graph and CoTs due to their high quality generation.
+All the prompt templates are given in Appendix C.
+Figure 4: Performance comparison between different strategies of CoT generation on TGQA.
+Strategy ER-T1 ER-T2 ER-T3 ER-T4
+ICL 0.13 0.13 0.31 0.10
+SFT 0.04 0.17 0.13 0.10
+SFT + bs 0.06 0.13 0.03 0.08
+SFT + bs + aug 0.03 0.05 0.04 0.05
+Table 3: Human evaluation on the CoTs generated by different strategies.
+T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error.
+(See Appendix B for error type explanations)
+  Model TGQA TimeQA TempReason
+Easy-mode Hard-mode OBQA-L2 OBQA-L3
+EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc
+${}^{{\dagger}}$ 0.410 0.608 - 0.600 0.682 - 0.556 0.641 - 0.260 0.450 - 0.238 0.418 -
+${}^{{\dagger}}$ 0.548 0.713 - 0.631 0.716 - 0.595 0.681 - 0.327 0.509 - 0.288 0.468 -
+${}^{{\dagger}}$ 0.640 0.778 - - - - - - - 0.318 0.496 - 0.261 0.430 -
+${}^{{\dagger}}$ 0.435 0.633 - 0.614 0.704 - 0.582 0.673 - 0.336 0.516 - 0.285 0.449 -
+${}^{{\dagger}}$ 0.461 0.660 - 0.637 0.723 - 0.605 0.693 - 0.374 0.549 - 0.334 0.493 -
+GPT-3.5 (ICL-SP) 0.598 0.699 - 0.640 0.668 - 0.512 0.506 - 0.303 0.409 - 0.365 0.453 -
+GPT-3.5 (ICL-CoT) 0.706 0.788 - 0.565 0.603 - 0.436 0.464 - 0.340 0.478 - 0.243 0.348 -
+$*{}^{*}$ (ICL-SP) 0.772 0.829 - 0.716 0.742 - 0.571 0.546 - 0.454 0.525 - 0.431 0.485 -
+$*{}^{*}$ (ICL-CoT) 0.821 0.865 - 0.662 0.693 - 0.618 0.636 - 0.388 0.480 - 0.352 0.447 -
+Llama2-7B (ICL-SP) 0.415 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354 0.205 0.344 0.226 0.044 0.084 0.153
+Llama2-7B (ICL-CoT) 0.548 0.686 0.578 0.367 0.425 0.393 0.302 0.354 0.360 0.233 0.410 0.263 0.179 0.357 0.187
+Llama2-13B (ICL-SP) 0.440 0.609 0.526 0.439 0.493 0.450 0.427 0.481 0.437 0.284 0.452 0.289 0.189 0.370 0.183
+Llama2-13B (ICL-CoT) 0.628 0.762 0.668 0.518 0.572 0.535 0.434 0.490 0.480 0.330 0.498 0.368 0.242 0.391 0.272
+Llama2-70B (ICL-SP) 0.618 0.736 0.665 0.583 0.627 0.631 0.493 0.537 0.551 0.358 0.491 0.387 0.128 0.181 0.148
+Llama2-70B (ICL-CoT) 0.761 0.838 0.796 0.552 0.612 0.623 0.447 0.501 0.512 0.325 0.477 0.345 0.303 0.393 0.318
+Llama2-13B (SFT-SP) 0.550 0.720 0.652 0.412 0.449 0.455 0.362 0.404 0.412 0.337 0.506 0.338 0.244 0.408 0.253
+Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379 0.256 0.433 0.281 0.285 0.409 0.305
+Llama2-13B (SFT-TGSR) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649 0.424 0.522 0.432 0.356 0.469 0.399
+Table 4: Main results using different models and strategies.
+We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc).
+Note: (1) Results with ${}^{{\dagger}}$ are reported in the original papers.
+We fine-tune and evaluate the models on our dataset.
+(2) Results with * are evaluated on 1000 random test samples.
+Figure 5: Performance comparison between different strategies on TimeQA and TempReason.
+To obtain a fair comparison, we use Llama2-13B as the base model for all strategies.
+The basic strategy is in-context learning with standard Input/Output (ICL-SP), which is used to calculate the performance changes.
+4.3 Main Results
+Can our strategy bring more reliable symbolic reasoning?
+Figure 4 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TGQA.
+It can be seen that LLMs learn better temporal reasoning with SFT than ICL.
+By providing CoTs with bootstrapping and our data augmentation strategies, the model performance gets further enhanced.
+Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies.
+Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error).
+It can be seen in Table 3 that our strategies reduce all types of error rates (ER).
+Can our two-step framework lead to better temporal reasoning performance?
+Table 4 shows the comparison between different models on all datasets.
+First, we observed that among all the LLMs with ICL, GPT-4 has the strongest performance as expected.
+For the Llama2 family, larger models have better performance due to advanced context understanding and improved generalization.
+We also found that CoTs not always bring better temporal reasoning due to unreliable intermediate results and hallucinations.
+More importantly, our model (SFT-TGSR), which is based on Llama2-13B, shows a comparable or even better performance than GPT-4 on all datasets.
+To prove the effectiveness of our framework, we also provide the results of some alternative strategies, where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively.
+It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets.
+We hypothesize that this improvement in performance is because our two-step neuro-symbolic process provides an easier path toward answering the temporal questions for LLMs.
+Does these learned capabilities of temporal reasoning generalize to other tasks?
+Figure 5 shows the comparison between different strategies on TimeQA and TempReason.
+In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning.
+More importantly, we observe that per-training on TGQA improves the model performance compared with ICL.
+Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions.
+Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the temporal reasoning capability of LLMs.
+4.4 Ablation Study
+We ablate different modules to see their contributions to the performance.
+Table 5 shows the performance comparison between different configurations.
+To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
+From the ablation study, we obtain some insights: (1) LLMs can benefit from explicitly presented (temporal) graph which is intuitive, concise and structured.
+(2) Given a reliable graph, symbolic reasoning brings better performance than free-text CoTs.
+(3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning.
+(4) The introduction of external knowledge such as mathematics and commonsense can augment the generation.
+5 Related Work
+Language-based Temporal Reasoning  Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023).
+The vision here is to help LMs understand temporal concepts.
+Existing methods mainly solve this problem with time-aware modeling.
+For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust TR.
+(Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge.
+Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
+Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023).
+Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods.
+For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.
+(Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing final answers.
+On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
+Neuro-symbolic Reasoning Recent research has shown the advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Hamilton et al., 2022).
+In particular, knowledge graphs (KGs) as the foundation of (temporal) symbolic reasoning has been widely adopted in the past (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021).
+Moreover,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs.
+In this paper, we build the connection between language-based and temporal KG-based temporal symbolic reasoning.
+This connection brings the potential for extending these KG-based methods to language-based tasks.
+Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain datasets, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to symbolic reasoning.
+Config. EM F1 Acc
+Baseline (SFT-CoT) 0.646 0.722 0.705
+TG 0.675 0.754 0.711
+TG + SR(bs) 0.723 0.797 0.736
+TG + SR(bs + aug) 0.782 0.838 0.795
+TG + EK + SR(bs) 0.753 0.803 0.776
+TG + EK + SR(bs + aug) 0.797 0.850 0.819
+Table 5: Ablation study results on TGQA.
+To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
+TG: temporal graph, SR: symbolic reasoning, EK: external knwoledge, bs: bootstrapping, aug: data augmentation.
+6 Conclusion
+TG-LLM, a new framework for language models, has been proposed to improve their performance on temporal reasoning.
+Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning steps.
+Experiments indicate that TG-LLM achieves better performance than existing methods.
+An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning.
+Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks.
+Limitations
+Graph-augmented approaches such as TG-LLM help language models better learn related concepts from the perspective of graph.
+Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021).
+Explicit in-context integration of commonsense presents opportunities for this task.
+Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data.
+There can be opportunities such as simulating an environment to provide feedback to LLMs  (Yao et al., 2023).
+For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range.
+In this way, we can further improve the quality of generation.
+Ethics Statement
+In this paper, we adopt YAGO11k for fine-tuning the language models.
+The dataset is publicly available, and is for research purposes only.
+We also use GPT model to generate text based on YAGO11k, for which OpenAI has been committed to addressing ethical considerations.
+In addition, we adopt TimeQA and TempReason for evaluation.
+Both datasets are publicly available, and are for research purposes only.
+However, they may still contain improper or harmful content.
+None of such content reflects the opinions of the authors.
+Appendix A TGQA Dataset Statistics
+In this section, we show statistics for our TGQA dataset.
+Figure 6 shows the number of facts distribution in training and test set, which determine the complexity of the task to some extent.
+Figure 7 shows the types of questions distribution in training and test set, where Q0: "Which event occurred first, <Event_A> or <Event_B>?", Q1: "Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in the chronological order?", Q2: "How long did the event <Event_A> last?", Q3: "", Q4: "How much time passed between the start of <Event_A> and the start of <Event_B>?", Q5: "What happened right before/after <Event_A> started?", Q6: "When did the <Event_A> occur?", Q7: "True or false: <Event_A> and <Event_B> happened at the same year?", Q8: "True or false: <Event_A> was still happening when <Event_B> started?".
+Note that we give eight categories of question in Table 2, since in a strict sense Q0 and Q1 can be considered as the same type.
+Among all the question types, Q1 has the largest portion since it has multiple variants.
+Similarly, Q7 with two variants also has a larger portion.
+To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores.
+We also provide examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.
+Entity Name Mapping
+Ori New
+Beverly_Adams Isabella_Thompson
+Edmonton Lancaster
+Al_Gore Chris_Evans
+Carrara_Stadium Maplewood_Arena
+Bend_Sinister_(novel) Lakefield_Chronicles_(novel)
+$\cdots$ $\cdots$
+Table 6: We use a global mapping for entity names from GPT-3.5 to avoid corpus leakage.
+External Knowledge
+1885 before 1893 before 1916 before 1918 before 1922
+before 1928 before 1941
+1918 - 1916 = 2
+1928 - 1893 = 35
+1928 - 1922 = 6
+1941 - 1918 = 23
+2 < 6 < 23 < 35
+Table 7: We integrate the necessary mathematics in context as external knowledge for temporal reasoning.
+Appendix B Experiment Details
+In this section, we present more experiment details for further research. We include data augmentation examples, model versions and evaluation tasks & metrics.
+Temporal Graph:
 […]
+[1] (John Thompson was born in Weston) starts at 1921;
 [2] (John Thompson run Pearl Network) starts at 1942;
@@
 run Pearl Network) ends at 1967.
-Graph-based QA (No Change):
+Graph-based QAs:
 […]
+Q: True or false: event (John Thompson owned Pearl
 Network) was longer in duration than event (Sophia
@@
 A: True.
-TempGraph (Inductive Transformation):
+Temporal Graph:
 […]
+[1] (James Brown was born in Oslo) starts at 1931;
 [2] (James Brown owned Iris Inn) starts at 1952;
@@
 Monaco) starts at 2005.
-Graph-based QA (Inductive Transformation):
+Graph-based QAs:
 […]
+Q: True or false: event (James Brown owned Iris Inn)
 was longer in duration than event (Ella Perry was
@@
 A: True.
-Table 3: We propose two special data augmentation strategies (adding disturbances and inductive transformation) for the neuro-symbolic reasoning.
-where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation name changing $f_{R}(\cdot)$.
-Inductive Transformation  In symbolic reasoning, a rule is considered as inductive if it is generalizable to different entities and times  (Xiong et al., 2022, 2023).
-For LLMs, this might be difficult to understand without enough training data since their fundamental principle is to learn the co-occurrence of tokens.
-To solve this problem, we introduce a new approach called inductive transformation.
-We globally map all the entity names in training data to some random names of the same type (e.g., "Knox Cunningham" $\rightarrow$ "Olivia Davis"; "Sydney" $\rightarrow$ "Cardiff"), and change the times based on a global time offset (e.g., the offset is "6 years": "June 1994" $\rightarrow$ "June 2000"; "2008" $\rightarrow$ "2014").
-For each relation, we generate these random names with LLMs by providing several examples of existing entity names.
-To avoid confusing the model, we only use obscure names that do not exist in our dataset.
-$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e)$ (6)
-$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c))$
-$a^{\prime\prime}=f_{T}(f_{E}(a))$
-where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity name changing and time changing, respectively.
-4 Experiments
-We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation?
-(2) Can our two-step framework lead to better temporal reasoning performance?
-(3) Can LLMs benefit from our data on other reasoning tasks?
-4.1 Experimental Setup
-In addition to TempGraph QA, we consider the existing dataset, TimeQA  (Chen et al., 2021), which is based on the texts from Wikipedia.
-It contains two difficulty levels: The easy-level split tends to be information extraction while the hard-level split involves temporal reasoning, i.e., understanding the relation between different temporal expressions.
-We evaluate our framework on both datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
-Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider the accuracy based on perplexity for LLM evaluations.
-Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format.
-This property actually hurts their performance under the commonly used metrics such as EM and F1.
-To solve this problem, we select the answers from a candidate set based on their perplexity from the LLM.
-For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them.
-For TempGraph QA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
-We compare our framework with leading LLMs including Llama2, GPT-3.5 and GPT-4.
-Llama2 Touvron et al.
-(2023b) is one of the most accepted open-source LLMs, while GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023) are two powerful closed-source models.
-We run Llama2 series on local machines (4-bit quantization is adopted for 70B version due to computational resource limitation), and use APIs provided by OpenAI for GPT series.
-We evaluate their few-shot ICL performance on the test set.
-To avoid LLMs generating free text which affects the evaluation with these metrics  (Yang et al., 2023a), we provide candidate answers for all questions under the ICL setting.
-The model versions and prompt templates are shown in Appendix B and C, respectively.
-4.2 Implementation Details
-We use Llama2-13B as the baseline due to computation resource limit.
-In experiments, we proved that our framework has comparable or even better temporal reasoning performance than one of the most powerful LLMs, GPT-4.
-We inject two adapters with selectors into the Llama2 model for the text-to-tempgraph translation and symbolic temporal reasoning.
-The adapters are trained in parallel.
-For inference, we first translate the original story into a temporal graph, and then perform symbolic temporal reasoning on it, i.e., the adapters are used in sequence.
-In addition, we use GPT-3.5 for the story creation & verification in TempGraph QA, entity and relation extraction, temporal info identification, graph construction & verification on other datasets, and symbolic CoTs bootstrapping.
-These examples used in ICL are created by GPT-4 with manual verification.
-We choose these models due to the high quality of their generation.
-All the prompt templates are shown in Appendix C.
-4.3 Main Results
-Can our strategies guide LLMs to perform more reliable symbolic reasoning?
-Figure 2 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TempGraph QA.
-To obtain a fair comparison, all the reasoning happens on the estimated graphs obtained from our text-to-graph translation module.
-It can be seen that LLMs learn better temporal reasoning with SFT than ICL.
-By providing CoTs with symbolic reasoning, the model performance gets further enhanced.
-There is an additional performance boost when using our data augmentation strategies.
-Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies.
-Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error).
-We measure the level of agreement among the three evaluators, and show the error rates from different types in Table 4.
-It can be seen that LLMs with ICL easily use either wrong information or erroneous external knowledge in reasoning.
-This issue can be alleviated with SFT, and we shows that using symbolic reasoning is better than free-text CoTs.
-Our data augmentation strategies further reduce all types of errors, and provides a promising direction for LLMs on reasoning tasks.
-Figure 2: Performance comparison between different strategies of CoT generation on TempGraph QA.
-Strategy ER-T1 ER-T2 ER-T3 ER-T4
-ICL 0.13 0.13 0.31 0.10
-SFT 0.04 0.17 0.13 0.10
-SFT + bs 0.06 0.13 0.03 0.08
-SFT + bs + aug 0.03 0.05 0.04 0.05
-Table 4: Human evaluation on the CoTs generated by different strategies. T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error.
-Model TempGraph QA TimeQA
-Easy-mode Hard-mode
-EM F1 Acc EM F1 Acc EM F1 Acc
-GPT-3.5 (ICL-SP) 0.559 0.699 - 0.640 0.668 - 0.512 0.506 -
-GPT-3.5 (ICL-CoT) 0.698 0.788 - 0.565 0.603 - 0.436 0.464 -
-GPT-4 (ICL-SP) 0.704 0.818 - 0.716 0.742 - 0.571 0.546 -
-$*{}^{*}$ (ICL-CoT) 0.798 0.833 - 0.662 0.693 - 0.618 0.636 -
-Llama2-7B (ICL-SP) 0.402 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354
-Llama2-7B (ICL-CoT) 0.548 0.686 0.578 0.367 0.425 0.393 0.302 0.354 0.360
-Llama2-13B (ICL-SP) 0.423 0.609 0.526 0.439 0.493 0.450 0.427 0.481 0.437
-Llama2-13B (ICL-CoT) 0.626 0.762 0.668 0.518 0.572 0.535 0.434 0.490 0.480
-Llama2-70B (ICL-SP) 0.578 0.736 0.665 0.583 0.627 0.631 0.493 0.537 0.551
-Llama2-70B (ICL-CoT) 0.760 0.838 0.796 0.552 0.612 0.623 0.447 0.501 0.512
-Llama2-13B (SFT-SP) 0.550 0.720 0.652 0.412 0.449 0.455 0.362 0.404 0.412
-Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379
-Llama2-13B (TempGraph) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649
-Table 5: Main results using different models and strategies.
-We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc).
-Results with * are evaluated on 1000 random test samples.
-Figure 3: Performance comparison between different strategies on TimeQA. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies.
-Can our two-step framework lead to better temporal reasoning performance?
-Table 5 shows the comparison between different models on both TempGraph QA and TimeQA.
-First, we observe that among all the LLMs with ICL, GPT-4 has the best temporal reasoning capability as expected.
-For Llama2 series, larger model has better performance due to advanced context understanding and improved generalization.
-We also find that CoTs not always bring better temporal reasoning.
-In both easy-mode and hard mode of TimeQA, CoTs hurt larger LMs such as GPT-3.5, GPT-4 and Llama2-70B due to unreliable intermediate results and hallucinations.
-More importantly, our model, which is based on Llama2-13B, shows a comparable or even better temporal reasoning performance than GPT-4 on both datasets.
-To prove the effectiveness of our framework, we also provide the results of some alternative strategies (ICL, SFT-SP, SFT-CoT), where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively.
-Compared with our framework, there are no text-to-graph translation and symbolic reasoning involved in SFT-SP and SFT-CoT.
-It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets.
-Compared with original text, we actually provide an easier path from temporal reasoning questions to their answers.
-Can LLMs benefit from our data on other reasoning tasks?
-Figure 3 shows the performance comparison between different strategies on TimeQA (ICL-SP, ICL-CoT, pre-training on TempGraph QA, SFT on TimeQA).
-It can be seen that the performance of in-context learning (either with standard I/O prompt or CoTs) is much lower than pre-training and SFT.
-In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning.
-More importantly, we observe that per-training on synthetic data improves the model performance compared with ICL, which proves the rationality of design and construction of our dataset.
-Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions.
-Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the reasoning capability of LLMs.
-4.4 Ablation Study
-We ablate different modules to see their contributions to the model performance.
-Table 6 shows the performance comparison between different configurations where TG, EK, SR denote temporal graph, external knowledge and symbolic reasoning, respectively.
-To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
-We use supervised fine-tuning on CoTs (SFT-CoT) as baseline.
-We improve the model performance 2.9% in EM, 3.2% in F1 and 0.6% in Acc by adding TG.
-Then we show that bootstrapping CoTs with symbolic reasoning brings further enhancement of 4.8% in EM, 4.3% in F1 and 2.5% in Acc.
-These is a significant boost in EM (5.9%), F1 (4.1%) and Acc (5.9%), when using our data augmentation strategies.
-Finally, the introduction of external knowledge exhibits meaningful results: it increases EM by 3%/1.5%, F1 by 0.6%/1.2%, and Acc by 4%/2.4% without/with data augmentation.
-From the ablation study, we give some insights for LLMs: (1) LLMs can benefit from (temporal) graph which is intuitive, concise and structured.
-(2) Given a reliable graph, symbolic reasoning brings better performance in reasoning than free-text CoTs.
-(3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning.
-(4) The introduction of external knowledge such as mathematics and commonsense can augment the generation of LLMs.
-Config EM F1 Acc
-Baseline (SFT-CoT) 0.646 0.722 0.705
-TG 0.675 0.754 0.711
-TG + SR(bs) 0.723 0.797 0.736
-TG + SR(bs + aug) 0.782 0.838 0.795
-TG + EK + SR(bs) 0.753 0.803 0.776
-TG + EK + SR(bs + aug) 0.797 0.850 0.819
-Table 6: Ablation Study Results on TempGraph QA. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
-5 Related Work
-Text-based Temporal Reasoning   Recently, text-based temporal reasoning has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023).
-The vision here is to help LMs understand temporal concepts, which are fundamental elements in human knowledge.
-Existing methods mainly solve this problem with time-aware modeling.
-For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust temporal reasoning.
-(Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge or extra temporal constraints.
-Although these methods make some progress, the underlying structure and logic of temporal reasoning are either ignored or not explicitly involved.
-Since LMs perform pre-training on human knowledge, the manual perspective of modelling might effectively improve their performance on this task.
-Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and their capability of reasoning (Valmeekam et al., 2022; Wei et al., 2022; Huang and Chang, 2023).
-Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modelling or hybrid methods.
-For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.
-(Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing the final answers.
-On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that requires reasoning.
-To the best of our knowledge, the exploration of the full capabilities of LLMs on temporal reasoning is still missing.
-Compared with other reasoning types, it is a form of integrated task that requires the knowledge of mathematics, logic and commonsense.
-Neuro-symbolic Reasoning   Recent research shows advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Susskind et al., 2021; Hamilton et al., 2022).
-Knowledge graph is an important representation as the foundation of symbolic reasoning.
-Existing works on symbolic reasoning over KGs include  (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021).
-On top of that,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs.
-In this paper, we build the connection between text-based and temporal KG-based temporal symbolic reasoning.
-This connection brings the potential to extend these KG-based methods on text-based tasks.
-Some examples include  (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023).
-Different from these methods, which limit their application to certain datasets, our framework has better generalization due to the text-to-graph translation part.
-6 Conclusion
-TempGraph-LLM, a new paradigm for language models, has been proposed to improve their performance on temporal reasoning.
-Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning process.
-Experiments indicate that TempGraph-LLM achieves better performance than existing methods.
-An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning.
-Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks with good problem formalization and methodology extension.
-Appendix A Dataset Statistics
-In this section, we show statistics for our TempGraph QA dataset.
-Figure 4 shows the number of facts in training and test set, which determine the complexity of task to some extent.
-Figure 5 shows the types of questions in training and test set.
-Different types of questions are exemplified in Table 1.
-Among all the types, question 1 has the largest portion since it has five variants.
-Similarly, question 7 with two variants also has a larger portion.
-To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores.
-Figure 4: Number of facts in TempGraph QA.
-Figure 5: Types of questions in TempGraph QA.
-Appendix B Model Versions
-All the model versions used in our experiments are listed below:
+Table 8: We propose special data augmentation strategies for symbolic reasoning over (temporal) graphs.
+Data Augmentation Examples  We show with an example our data augmentation strategies in Table 8.
+For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on Graph-based QAs.
+For TGs with entity names/times mapping, we also need to change with the corresponding entity names/times in Graph-based QAs to ensure the consistency.
+Model Versions  All the model versions of LLMs used in our experiments are listed below:
   • Llama2-7B (meta-llama/Llama-2-7b-hf)
@@
   • GPT-4 (gpt-4-1106-preview)
+For the Llama2 family, all the model weights can be downloaded from the platform of Hugging Face. For the GPT models, all the model weights can be accessed through the OpenAI APIs.
+Evaluation Tasks & Metrics  In addition to TGQA, we consider the existing datasets TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries.
+To be specific, TimeQA  (Chen et al., 2021) contains two difficulty levels (see Table 9).
+The easy-level split tends to be information extraction while the hard-level split involves understanding the relation between different temporal expressions.
+TempReason  (Tan et al., 2023a) contains three levels (L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation) and three settings (OBQA: open-book QA, CBQA: close-book QA, ReasonQA: facts-based QA) (see Table 10).
+We observe that some stories in TempReason are incomplete which partially leads to the low accuracy of LLMs.
+We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
+Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations.
+TimeQA
+George Washington (February 22, 1732 – December 14,
+1799) was an American Founding Father, military offic-
+er, politician and statesman who served as the first $\cdots$
+Questions (Easy-mode):
+What position did George Washington hold in June 17-
+75?
+What position was held by George Washington between
+1778 and 1788?
+$\cdots$
+Questions (Hard-mode):
+George Washington took which position before 1778?
+What was George Washington’s position in early 1780s?
+$\cdots$
+Table 9: Example questions of two difficulty levels in TimeQA.
+Easy-mode: the query time expression is explicitly mentioned in the story.
+Hard-mode: obtaining the answer needs inference based on the temporal relation between the query time expression and the one mentioned in the story.
+TempReason
+Lionel Andrés "Leo" Messi (born 24 June 1987) is an Ar-
+gentine professional footballer who plays as a forward for
+and captains both Major League Soccer club $\cdots$
+Questions (L1):
+What is the year after 2010?
+$\cdots$
+Questions (L2):
+What team did Leo Messi play for in 2010?
+$\cdots$
+Questions (L3):
+What team did Leo Messi play for after Barcelona?
+$\cdots$
+Table 10: Example questions of three difficulty levels under the OBQA setting in TempReason. L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation.
+Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format.
+This property actually hurts their performance under the commonly used metrics such as EM and F1.
+To solve this problem, we select from a candidate set the answer with the lowest perplexity from LLMs as their prediction.
+For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them.
+Since there are multiple categories of questions in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
+  Model TGQA
+Q0 Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Total
+${}^{\ddagger}$ (0-shot SP) 0.801 0.435 0.660 0.355 0.930 0.955 0.710 0.315 0.617 0.642
+GPT-3.5 (1-shot SP) 0.842 0.411 0.656 0.555 0.924 0.650 0.500 0.353 0.495 0.598
+${}^{\ddagger}$ (0-shot CoT) 0.930 0.565 0.680 0.665 0.920 0.945 0.695 0.395 0.596 0.710
+GPT-3.5 (1-shot CoT) 0.884 0.457 0.783 0.791 0.921 0.832 0.755 0.357 0.574 0.706
+${}^{{\ddagger}*}$ (0-shot SP) 0.801 0.689 0.920 0.620 0.930 0.940 0.758 0.575 0.737 0.776
+$*{}^{*}$ (1-shot SP) 0.820 0.660 0.870 0.630 0.910 0.990 0.760 0.610 0.700 0.772
+${}^{{\ddagger}*}$ (0-shot CoT) 0.950 0.692 0.901 0.881 0.919 0.980 0.778 0.647 0.858 0.848
+$*{}^{*}$ (1-shot CoT) 0.930 0.758 0.903 0.869 0.892 0.920 0.758 0.667 0.680 0.821
+Llama2-7B (1-shot SP) 0.554 0.216 0.163 0.147 0.353 0.924 0.683 0.115 0.579 0.415
+Llama2-7B (1-shot CoT) 0.660 0.264 0.484 0.603 0.693 0.947 0.632 0.079 0.574 0.548
+Llama2-13B (1-shot SP) 0.442 0.341 0.353 0.192 0.686 0.640 0.561 0.163 0.580 0.440
+Llama2-13B (1-shot CoT) 0.597 0.263 0.670 0.678 0.881 0.947 0.736 0.258 0.622 0.628
+Llama2-70B (1-shot SP) 0.752 0.447 0.520 0.349 0.941 0.891 0.665 0.413 0.585 0.618
+Llama2-70B (1-shot CoT) 0.980 0.490 0.878 0.832 0.944 0.878 0.807 0.294 0.750 0.761
+Llama2-13B (1-shot SFT - TG) 0.878 0.515 0.656 0.866 0.835 0.845 0.726 0.369 0.383 0.675
+Llama2-13B (1-shot SFT - TG + SR(bs)) 0.947 0.732 0.747 0.825 0.931 0.987 0.495 0.466 0.649 0.753
+Llama2-13B (1-shot SFT - TG + SR(bs + aug)) 0.931 0.710 0.729 0.860 0.927 0.977 0.627 0.534 0.739 0.782
+Llama2-13B (1-shot SFT - TG + EK + SR(bs + aug)) 0.944 0.775 0.729 0.849 0.924 0.980 0.783 0.506 0.681 0.797
+[…]
+We explain the categories of errors in human evaluation on CoT generation here (see Table 3).
+T1 means that LLM uses wrong information such as wrong start/end time during reasoning.
+T2 means that LLM makes logical errors, e.g., after determining "Event A ends before Event B starts", considering the statement as true that "Event A was still happening when Event B starts".
+T3 means that LLM makes errors on external knowledge, e.g., claiming that "the date 1978 is after 1983".
+T4 means that there exists error in the extracted TG which leads to the wrong conclusion.
 Appendix C Example Prompts
 In this section, we show example prompts used in our framework.
-To be specific, Table 7 shows an example of the graph-based story generation in TempGraph QA; Table 8 presents an example of the automatic story-tempgraph alignment verification in TempGraph QA; Table 9 provides an example of the temporal info identification in TimeQA; Table 10 illustrates an example of the graph construction in TimeQA; Table 11 gives an example of the automatic tempgraph-QA alignment verification in TimeQA; Table 12 depicts an illustrative case of the CoTs bootstrapping in TempGraph QA.
+To be specific, Table 12 shows an example of the graph-based story generation in TGQA; Table 13 presents an example of the automatic story-temporal graph alignment verification in TGQA; Table 14 provides an example of the temporal info identification in TimeQA; Table 15 illustrates an example of the graph construction in TimeQA; Table 16 gives an example of the automatic temporal graph-QA alignment verification in TimeQA; Table 17 depicts an illustrative case of the CoTs bootstrapping in TGQA.
+Appendix D Fine-grained Results
+We provide fine-grained results on TGQA in Table 11.
+In TGQA, there are nine types of questions as explained in Appendix A.
+We show all the models with different strategies here.
+Zero-shot performance of LLMs is also considered to investigate the effect of in-context examples.
+Note that we do not include zero-shot performance of the Llama2 family since it is not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results.
+For zero-shot learning results, the format of generation is not guaranteed.
+The original accuracy is very low since rule-based parser cannot handle it.
+Instead, we use GPT-3.5 as parser to extract the answer from the generation.
+Also, we only adopt one example in context due to context length limitation.
+It can be seen that in-context examples are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4.
+Undoubtedly, GPT-4 obtains the best in-context learning performance on most categories.
+However, for some categories, the best performance of GPT-4 is around 0.6 or 0.7, which is not sufficiently well.
+On the other hand, we show that our strategies improve the performance step by step.
+Note that our model achieves the best overall performance by improving all the categories.
+To be specific, symbolic reasoning brings better performance for most categories.
+Furthermore, data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.
 (Molly Adams was born in Seattle) starts at 1896
@@
 Based on the above facts, generate a short story for me.
-Table 7: Example of the graph-based story generation in TempGraph QA.
+Table 12: Example of the graph-based story generation in TGQA.
 Once upon a time, in the vibrant city of Seattle, two remarkable individuals, Molly
@@
 When did the event (Molly Adams was married to Liam Thomas Dawson) end?
-Table 8: Example of the automatic story-tempgraph alignment verification in TempGraph QA.
+Table 13: Example of the automatic story-temporal graph alignment verification in TGQA.
 Knox Cunningham
@@
 Which position did Knox Cunningham hold from May 1955 to Apr 1956?
-Table 11: Example of the automatic tempgraph-QA alignment verification in TimeQA.
+Table 16: Example of the automatic temporal graph-QA alignment verification in TimeQA.
 Example:
@@
 Answer: Let’s think step by step.
-Table 12: Example of the CoTs bootstrapping in TempGraph QA.
+Table 17: Example of the CoTs bootstrapping in TGQA.

### v3

@@
 Recent studies have introduced various methods to mitigate these limitations.
-Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate contextual details.
-In this paper, we propose TG-LLM, a new framework towards language-based TR.
-To be specific, we first teach LLM to translate the context into a temporal graph (TG).
-A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for fine-tuning on this graph translation task.
-We confirm in experiments that the capability of TG extraction learned on our dataset can be transferred to other TR tasks and benchmarks.
-On top of that, we guide LLM to perform symbolic reasoning over the TG via Chain of Thoughts (CoTs) bootstrapping and special data augmentation strategies.
-We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free-form text.
+Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate temporal logic.
+In this paper, we propose TG-LLM, a novel framework towards language-based TR.
+Instead of reasoning over the original context, we adopt a latent representation, temporal graph (TG) that facilitates the TR learning.
+A synthetic dataset (TGQA), which is fully controllable and requires minimal supervision, is constructed for fine-tuning LLMs on this text-to-TG translation task.
+We confirmed in experiments that the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks.
+On top of that, we teach LLM to perform deliberate reasoning over the TGs via Chain of Thought (CoT) bootstrapping and graph data augmentation.
+We observed that those strategies, which maintain a balance between usefulness and diversity, bring more reliable CoTs and final results than the vanilla CoT distillation.
+[^1] The code is available at https://github.com/xiongsiheng/TG-LLM.
 1 Introduction
 As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception.
-It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations.
-Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023).
-However, they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023).
-TR is more challenging since it requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
-Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and neuro-symbolic reasoning.
-Recent works mainly adopt the general reasoning methods to investigate and improve the TR capability of LLMs.
-For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different TR tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).
+It is not just about understanding basic concepts such as ordering or duration; it extends to more intricate aspects, e.g., task planning or causal relation discovery.
+Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have emerged with some reasoning capabilities  (Huang and Chang, 2023).
+However, there is observation that they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023), preventing their applications from solving complex real-world problems.
+In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
+Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and temporal graph reasoning.
+Recent works mainly adopt general approaches to investigate and improve the TR capability of LLMs.
+For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2023a).
+Similarly,  (Wei et al., 2023b) designs several specific types of prompts as prompt tuning.
 (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT).
-Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL).
 In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
 Despite the effectiveness of such methods, they either ignore or not explicitly involve the intrinsic nature of TR.
-Human performs complex TR on a timeline of events which are aligned with the entities and relations.
+Humans perform complex TR on a timeline of events which are aligned with the entities and relations.
 These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information.
-Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline of events from the text, and then perform deliberate reasoning on it.
-However, the high-quality extraction of the aligned timeline is a challenging task itself due to the lack of ground truth.
-To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision.
-We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks and benchmarks.
-Given a reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on it: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR?
-(2) How can one guide LLMs to perform rigorous reasoning?
-Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps).
-As for (2), there are two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al.
-(2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023).
+In other words, the aligned timeline (more generally, the temporal graph, TG) serves as a latent representation to help humans understand the patterns in TR.
+However, due to the lack of ground truth, the high-quality TG translation is a challenging task for most TR benchmarks.
+To solve this problem, we propose a synthetic dataset (TGQA), which is fully controllable and requires minimal supervision.
+We demonstrate the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks.
+Given a reliable TG, the key challenges of teaching LLMs TR include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR?
+Prior works  (Lewis et al., 2021) show that explicitly introducing knowledge in context enhances the performance of LLMs.
+In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps).
+(2) How can one teach LLM to perform deliberate reasoning?
+Generally, there exist two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al.
+(2023); (ii) using LLMs directly as the reasoning engine  (Zhu et al., 2023).
 For (i), the difficulty lies in accurate translation  Yang et al.
 (2023c) and the limited expressive power of formal logic.
-For (ii), there is no guarantee for the correctness of the generated text especially when training data are insufficient  (Yang et al., 2023b).
-To solve these problems, we propose a symbolic-integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning.
-Furthermore, we design special data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
+For (ii), there is no guarantee for the correctness of generated intermediate steps especially with insufficient training data  (Yang et al., 2023b).
+We adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning.
+We further improve the model performance with graph data augmentation which mitigates the data deficiency in TR tasks.
 To be specific, our contributions are summarized as follows:
   • We propose a new paradigm, TG-LLM, for language-based TR.
-In this framework, we first translate the context into a temporal knowledge graph, and then perform symbolic reasoning on it.
-Extensive experiments prove that our novel approach results in superior performance compared to the conventional approach.
-  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLM to utilize symbolic reasoning, which brings better performance than free-form text.
-  • We present a pipeline to create a synthetic dataset for question answering that requires TR.
+In this framework, we first translate the context into a latent representation (temporal graph), and then perform reasoning on it.
+Extensive experiments prove that our novel approach results in superior performance compared to the baselines.
+  • We design two approaches including Chain of Thought bootstrapping and graph data augmentation to teach LLM to generate consistent and faithful CoTs, which brings better performance than the vanilla CoT distillation.
+  • We present a pipeline to create a synthetic dataset (TGQA) for question answering that requires TR.
 It is fully controllable and requires minimal supervision for text-temporal graph alignment.
-We show in experiments that fine-tuning on our dataset benefits LLMs on other TR tasks and benchmarks.
+We show in experiments that fine-tuning on our dataset benefits LLM on other TR tasks and benchmarks.
 Temporal Graph (sub; rel; obj; start/end; time):
@@
 Table 1: Each sample from TGQA dataset is in the form of (story, temporal graph, questions, answers).
-2 Dataset Construction
 Reasoning Type Question Answer
@@
 Temporal How much time passed between the start of <Event_A> and the start of <Gap_between_Event_A
-Relation <Event_B>? _and_B_start_time>
+Relation <Event_B>? _and_Event_B_startTime>
 What happened right before/after <Event_A> started? <Event_B> right before/
 after <Event_A>
-Facts When did the <Event_A> occur? <Event_A_start_time>
+Facts When did the <Event_A> occur? <Event_A_startTime>
 Extraction
@@
 Analysis
-Table 2: Reasoning types with corresponding questions and answers in TGQA.
-We propose a pipeline for the construction of our synthetic dataset called TGQA.
-Each sample, as shown in Table 1, is in the form of (story, temporal graph, questions, answers).
-The temporal graph (TG) is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), with a restriction on the number of events.
-Each story is translated from the TG by GPT-3.5.
-Finally, by using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph.
-We obtain 500 samples for training/validation and another 100 for test, with about 30 QA pairs in a single sample.
-Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of temporal questions as shown in Table 2.
-Below we provide the details of each step in the dataset construction pipeline.
-Step 1: Graph Splitting & Masking  Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them into text all at once.
-Given a specific entity, we extract its third-order neighbors and all the events happening between them.
-Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training/validation and test samples.
-In addition, we consider the noticeable corpus leakage problem of LLMs, where prior knowledge of test data has been implicitly obtained from the extensive pre-training process.
-Thus, a masking strategy, i.e., changing entity names into random ones of the same type, is adopted.
-We generate a global mapping for entity names with GPT-3.5 (see Table 6).
-To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
-Step 2: Graph-based Open QA Creation  Based on the given TG, we generate a story and multiple QAs.
-We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts.
+Table 2: Reasoning types with the corresponding questions and answers in TGQA.
+2 Dataset Construction
+In this section, we present the construction pipeline for TGQA dataset that is fully controllable and requires minimal supervision for text-temporal graph alignment.
+Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of TR questions (Table 2), and the ground-truth timeline.
+We first split the large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), into subgraphs with a restriction on the number of events, and anonymize the entities to avoid data leakage.
+Then each story is translated from the subgraph by GPT-3.5  (Ouyang et al., 2022).
+By using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph.
+Finally, to reduce the noise introduced from the misalignment between the subgraph and generated story, we propose a semi-automatic verification method.
+Step 1: Graph Splitting & Anonymization.
+Existing temporal knowledge graphs (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) usually have a large size.
+To facilitate the learning process, we split YAGO11k into subgraphs for story generation.
+Specifically, given a certain entity, we find its neighbors within three hops, and extract all the events happening between them.
+Since LLMs are supposed to do reasoning instead of memorization, we ensure that no overlapping exists between the events in training, validation and test sets.
+Additionally, we notice the data leakage problem of LLMs, i.e., prior knowledge of the test data has been implicitly obtained from the extensive pre-training.
+Thus, an anonymization strategy, i.e., changing entity names into random ones of the same type, is adopted.
+We generate a global mapping for entity names with GPT-3.5.
+To avoid confusion, we adopt obscure names that do not exist in YAGO11k.
+Step 2: Graph-based Open QA Creation.
+In TGQA, each sample is in the form of (temporal graph, story, questions, answers) (Table 1).
+Based on the given subgraph, we generate a story and multiple QAs.
+We first ask GPT-3.5 to write a story based on the subgraph with the requirements to include all the events.
 It is observed that GPT-3.5 tends to ignore the end time of some events in the created story.
-To solve this problem, we split the starting and ending of a persistent event into two transient ones (see Table 1).
-To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction.
-Given these categories, we design multiple diverse question types (see Table 2), and use a rule-based Python script to generate the corresponding Qs and As.
-Step 3: Quality Control  Although LLMs are good at story generation, there is a risk of misalignment between the given TG and the generated text.
-To avoid the interference caused by misalignment, we propose a semi-automatic way for verification.
-Given a generated story, we ask LLMs the time of each event in the graph.
-If LLMs cannot give the correct answer, we consider the event as possibly missing in the story.
-Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors and only manually inspect the set of unanswered questions by LLM.
-Note that only the quality of story-TG alignment needs manual inspection, since all the QAs are generated from the TG with specific rules.
+To solve this problem, we separate the start and end time of the same event in prompt.
+On the other hand, to obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction.
+Given these categories, we design multiple question types (Table 2) with a rule-based Python script to generate the corresponding Qs and As.
+Step 3: Quality Control.
+In TGQA, we might introduce noise from the misalignment between the given subgraph and generated story.
+To reduce the noise, we propose a semi-automatic verification method.
+Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors, and only manually inspect the set of unanswered questions by LLM.
+Specifically, given a generated story, GPT-3.5 is queried on the time of each event in the graph.
+If it cannot give the correct answer, we consider the event as possibly missing in the story, which requires manual verification.
+We proved the effectiveness of our semi-automatic pipeline with fully manual verification of the test stories.
+Note that supervision is only required for story-TG alignment verification, since all the QAs are generated from rules.
 We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
 3 TG-LLM
-Motivated by human perception, we propose a new paradigm called TG-LLM. We first translate the text into a temporal graph, and then guide the LLM to perform symbolic reasoning on it.
+Motivated by human perception, we propose a new paradigm called TG-LLM. We first translate the text into a temporal graph (TG), and then guide the LLM to perform deliberate reasoning on it.
 3.1 Text-to-TG Translation
-TG serves as the foundation of the following symbolic reasoning process.
-Although LLMs can extract fact/events from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs, i.e, LLMs making mistakes or using different entity/relation names or time expressions.
-Thus, we provide a deliberate pipeline for high-quality TG translation.
-Ground-truth TG Generation   For datasets such as TGQA, each story is corresponding to a verified TG.
-Thus, we can directly fine-tune LLMs on this task with the ground truth.
-However, the biggest challenge for most real-world applications is the lack of ground truth.
-Instead of using the naive estimation from LLMs, we provide a pipeline for high-quality generation.
-(1) Entity & Relation Extraction: To align the entity/relation names in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context.
-To obtain the entity/relation names in pre-defined QAs, we consider two ways: parsing with rules or using LLMs.
-(2) Temporal Info Identification: To identify all the time expressions, LLMs are first used to extract them from the context, and then a rule-based Python script is adopted for time normalization and invalid output filtering.
-For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, which are introduced as external knowledge in temporal reasoning (see Table 7).
-(3) Graph Construction & Verification: We generate the graph using LLMs with ICL.
-To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (see Table 15).
-For those without explicit expressions, we ask LLMs to provide the temporal order of different events.
-We further semi-automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph.
-For those graphs where LLMs give the wrong answers, we manually check and correct them.
-Figure 2: In chain of thoughts (CoTs) bootstrapping, we accept CoTs leading to correct answers and sample them according to their contrastive learning scores.
-Supervised Fine-tuning  We found out in our experiments that providing a chronological list of events helps LLMs provide better temporal reasoning.
-In addition, as shown in Table 1, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events.
-With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times.
+Although LLM (with ICL) might have such capability, we observed a misalignment between the generated TG and pre-defined QAs, i.e, LLM making mistakes or focusing on irrelevant events.
+Since TG serves as the foundation of the following deliberate reasoning process, we provide a pipeline for high-quality TG data generation to fine-tune the LLM.
+Ground-truth TG Generation.
+For some datasets such as TGQA, a verified TG (corresponding to the story) is provided.
+We can directly fine-tune the LLM on this task with the ground truth.
+However, for most real-world applications, the biggest challenge is the lack of ground truth TG.
+We provide a pipeline for high-quality TG data generation.
+We first extract all the entities and relations from the QAs to help LLM focus on specific events.
+We then identify all the valid time expressions in the story.
+We provide these information to LLM for better alignment in TG construction, and verify the generated TG in a semi-automatic way.
+(1) Entity & Relation Extraction: To obtain the entity and relation in pre-defined QAs, we consider two ways: parsing with rules or using GPT-3.5.
+Rule-based parsing, if applicable, is more efficient and reliable, which is prioritized in our experiments.
+(2) Temporal Info Identification: To identify all the valid time expressions, we first use GPT-3.5 to extract from the story, and then adopt a rule-based Python script for invalid output filtering and normalization.
+Besides bringing better alignment, the pre-processing of time expressions facilitates the introduction of external knowledge (their temporal relations and time gaps) into TR.
+We proved in experiments that explicitly introducing this information further improves model performance.
+(3) TG Construction & Verification: We generate the TG using GPT-3.5 with ICL.
+Specifically, we provide several in-context demonstrations which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (Table 15).
+For those stories without explicit time expressions, we skip Step (2), and ask GPT-3.5 to provide the temporal order of events.
+Similar to TGQA, we verify the generated TG by first asking GPT-3.5 the pre-defined QAs, and then manually checking if GPT-3.5 fails.
+Figure 2: In Chain of Thoughts (CoTs) bootstrapping, we accept CoTs leading to correct final answers and sample them according to their contrastive learning scores.
+Supervised Fine-tuning.
+We first conducted experiments to decide the best in-context format of TG.
+It is found out that providing a chronological list of events helps LLM perform better TR.
+In addition, it offers advantages to separate the start and end time of the same event.
+With these steps, the TG in context is transferred into a timeline with the alignment between entities, relations and times.
 We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021).
@@
 In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
-3.2 Neuro-symbolic Temporal Reasoning
-Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by CoTs bootstrapping and data augmentation.
+3.2 Temporal Graph Reasoning
+Given the generated TGs, we teach LLM deliberate reasoning with SFT enhanced by CoT bootstrapping and graph data augmentation.
 3.2.1 Bootstrapping Chain of Thoughts
-It is observed that SFTs on reliable CoTs bring better performance in reasoning tasks than those on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023).
-However, it is challenging to obtain reliable CoTs from LLMs for symbolic reasoning tasks  (Saparov and He, 2023).
-Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples.
-To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{M}\}$.
-For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask GPT-3.5/GPT-4 to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$.
-These CoTs will be manually verified and used as few-shot ICL examples.
-When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})$ from GPT-3.5, and only accept the one leading to correct answers, i.e., $\hat{a}_{j^{\prime},k}=a^{*}_{j^{\prime}}$.
-In addition, inspired by  (Wang et al., 2023), we compare different CoTs using contrastive learning given in Eq.
-2.
-The core idea is to choose CoTs that make the correct answer more plausible than other candidate answers.
-For the generated CoTs, we use weighted sampling according to the score-based probability given in Eq.
-1.
+It is observed that SFT on reliable CoTs brings better reasoning performance than that on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023).
+Since asking humans to create the CoT data is not scalable, we use LLM to replace humans.
+This task is non-trivial for reasoning over knowledge graphs  (Saparov and He, 2023).
+In this section, we propose a bootstrapping pipeline, i.e., given a query, using LLM to generate several CoTs and selecting them as training data with a weighted sampling strategy.
+Compared with the conventional Best-of-N sampling, our proposal allows more training data diversity.
+Specifically, we first prepare ICL examples for high-quality CoT generation.
+To facilitate learning, the pre-defined QAs are classified into different categories $\{Q_{1},Q_{2},\cdots$, $Q_{M}\}$.
+For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$, where $g$, $e$, $q$, $a$ denote TG, external knowledge, question and answer, respectively, and ask both GPT-3.5 and GPT-4 to provide diverse CoTs $\{c_{j}\}^{N_{i}}_{j=1}$.
+These CoTs will then be manually verified as ICL examples.
+Given a new training sample $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}})$, we bootstrap $K$ CoTs with final answers $\{(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})\}^{K}_{k=1}$ from GPT-3.5.
+We first refuse the CoTs leading to incorrect answers, i.e., $\hat{a}_{j^{\prime},k}\neq a^{*}_{j^{\prime}}$, where $a^{*}_{j^{\prime}}$ denotes the correct answer.
+For the accepted CoTs, we consider a weighted sampling strategy to balance usefulness and diversity.
+The sampling probability $P_{sample}(\cdot)$ is based on a contrastive learning $score$ (Eq.
+1).
+The $score$ design, inspired by  (Wang et al., 2023), considers both perplexity $P(\cdot)$ of the correct answer and plausibility growth $G(\cdot)$ (Eq.
+2).
+The definition of $G(\cdot)$ and $P(\cdot)$ are given in Eq.
+3 and Eq.
+4, respectively.
 $P_{sample}(c_{k})=\operatorname{softmax}(score(c_{k}))$ (1)
@@
 $P\left(a\mid g,e,q,c_{k}\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c_{k},t_{<l}\right)$ (4)
-where $P_{sample}(\cdot)$ is the sampling probability for CoTs selection, $score(\cdot)$ is the contrastive learning score for each CoT, $g,e,q,c$ denote the TG, external knowledge, question, and CoT, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
-Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method has more diversity and does not require prior knowledge for the reasoning process.
-3.2.2 Data Augmentation
-Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023).
-To address the issue, we propose special strategies for data augmentation.
-Essentially, our framework involves two steps: text-to-TG translation and symbolic temporal reasoning.
-At the test stage, reasoning happens on the noisy estimated graphs, which impacts the overall performance.
-Specifically, for the reasoning module, we train with ground-truth/verified graphs while in the test time, the estimated graph is used for symbolic reasoning.
-To improve the robustness of reasoning, we introduce some disturbances on the graphs during training.
-Note that the type of disturbances should be carefully designed in order to avoid confusing the model.
-We first investigate two types of disturbances: (i) remove irrelevant edges and (ii) replace edges by changing relation names, given in Eq.
-5 and Eq.
-6, respectively.
-An edge is considered irrelevant if it is not involved in both QA and CoT.
+where $c_{k}$ denotes the current CoT in consideration; $a^{\prime}$ denotes a certain wrong answer from the candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $t_{l}$ denotes the $l$-th token in $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
+3.2.2 Graph Data Augmentation
+Compared with other tasks, reasoning suffers more from data insufficiency since more information (such as evidence, arguments, and logics) are involved in the intermediate steps  (Huang and Chang, 2023).
+To address this issue, we propose several graph data augmentation strategies (Figure 3).
+Our framework involves two steps: text-to-TG translation and temporal knowledge graph reasoning.
+Notably, for the reasoning part, we train with ground-truth/verified TGs but infer on the estimated graphs.
+This discrepancy hurts the robustness of reasoning.
+Thus, we introduce some disturbances on the TGs during training.
+Note that the type of disturbances should be carefully designed in order to avoid confusing the LLM.
+We first investigate two types of disturbances: (i) remove irrelevant edges (Eq.
+5) and (ii) replace edges by using relation synonyms (Eq.
+6).
+An edge (event) is considered irrelevant if not involved in both QA and CoT.
 $F_{irr}\in\{F:P(c|g_{\setminus\{F\}},e,q)\approx P_{0}\}_{F\in g}$ (5)
 $F^{\prime}\in\{f_{R}(F):P(c|f_{R}(g),e,q)\approx P_{0}\}_{F\in g}$ (6)
-Figure 3: We have several strategies for data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
+Figure 3: We have several strategies for graph data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
 where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation names $f_{R}(\cdot)$.
-Furthermore, to make sure LLM learns the underlying principle of symbolic temporal reasoning instead of just memorizing the training data, we introduce another two types of disturbances: (iii) globally map all the entity names in training data to some random names of the same type , and (iv) change the times based on a global offset, given in Eq.
-7.
-For each relation, we generate these random names with GPT-3.5 by providing several examples of existing names.
-$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e),$ (7)
-$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c)),$
-$a^{\prime\prime}=f_{T}(f_{E}(a))$
+Furthermore, to make sure LLM learns the underlying logic of TR instead of just memorizing semantic information, we introduce another two types of disturbances: (iii) globally map all the entity names in training data to some random names of the same type (Eq.7) , and (iv) change the times based on a global offset (Eq.7).
+For each relation, we generate these random entity names with GPT-3.5 by providing several examples of existing names.
+$g^{\prime}=f_{T}(f_{E}(g)),\ e^{\prime}=f_{T}(e),$ (7)
+$q^{\prime}=f_{T}(f_{E}(q)),\ c^{\prime}=f_{T}(f_{E}(c)),$
+$a^{\prime}=f_{T}(f_{E}(a))$
 where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity names and time changing, respectively.
 4 Experiments
-We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation?
-(2) Can our two-step framework lead to better temporal reasoning performance?
-(3) Does these learned capabilities of temporal reasoning generalize to other tasks?
+We aim to answer the following research questions in our experiments: (1) Can our strategies (CoT bootstrapping and graph data augmentation) bring more reliable reasoning over TGs?
+(2) Can our two-step framework lead to better TR performance?
+(3) Do these learned capabilities of TR generalize to other tasks?
 4.1 Experimental Setup
-In addition to TGQA, we consider the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries.
-Compared with these datasets, TGQA has a better quality due to the controlled generation, and more diverse question categories (see examples from other datasets in Appendix B).
-We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
-Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations, i.e., selecting from a candidate set the answer with the lowest perplexity from LLMs as their prediction.
-(See rationales for evaluation metrics selection in Appendix B)
-We compare our framework [^1] Code and data available at https://github.com/xiongsi heng/TG-LLM.
-with leading LLMs including Llama2 Touvron et al.
+We demonstrate TG-LLM is a general framework by applying it to, besides TGQA, the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries.
+Examples from other datasets are listed in Appendix B.
+We thoroughly evaluate our framework on all the datasets with a combination of the metrics, i.e., token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
+Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluation, i.e., selecting from a candidate set the final answer with the lowest perplexity as the prediction.
+The construction of the candidate set for all datasets and detailed explanations are listed in Appendix B.
+We primarily compare our framework with the leading LLMs, i.e., Llama2 Touvron et al.
 (2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023).
-We run the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and use APIs provided by OpenAI for GPT models.
-We evaluate their few-shot ICL performance on the test set.
-To avoid LLMs generating free text which affects the evaluation, we provide candidate answers for all questions.
+Specifically, we ran the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and used APIs provided by OpenAI for GPT models.
+We evaluated their few-shot ICL performance on the test set.
+To prove the effectiveness of our method, we also show the model performance with SFT on standard Input/Output prompts (SP) and CoTs from GPT-3.5.
 The model versions and prompt templates are provided in Appendix B and C, respectively.
@@
 We use Llama2-13B as the baseline due to limited computational resources.
-We inject two adapters with selectors into the model for the text-to-TG translation and symbolic temporal reasoning.
+We inject two adapters with selectors into the base model for the text-to-TG translation and temporal knowledge graph reasoning.
 The adapters are trained in parallel.
-For inference, we first translate the original story into a temporal graph, and then perform symbolic reasoning on it, i.e., the adapters are used in sequence.
-In addition, we use GPT-3.5/GPT-4 for all the generation of story, graph and CoTs due to their high quality generation.
+For inference, we first translate the original story into a temporal graph, and then perform reasoning on it, i.e., the adapters are used in sequence.
+For data generation, we use GPT-3.5 for story, TG and CoT generation, and the verification of stories and TGs.
+We use GPT-4 to create the ICL demonstrations of CoT generation, due to its high generation quality.
 All the prompt templates are given in Appendix C.
-Figure 4: Performance comparison between different strategies of CoT generation on TGQA.
+Figure 4: Performance comparison between different CoT generation strategies on TGQA.
 Strategy ER-T1 ER-T2 ER-T3 ER-T4
@@
 SFT + bs + aug 0.03 0.05 0.04 0.05
-Table 3: Human evaluation on the CoTs generated by different strategies.
-T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error.
-(See Appendix B for error type explanations)
+Table 3: Human evaluation on the generated CoTs by different strategies.
+ER: error rate; T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error.
+(Error type explanations are listed in Appendix B.)
     Model TGQA TimeQA TempReason
@@
 EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc
-${}^{{\dagger}}$ 0.410 0.608 - 0.600 0.682 - 0.556 0.641 - 0.260 0.450 - 0.238 0.418 -
-${}^{{\dagger}}$ 0.548 0.713 - 0.631 0.716 - 0.595 0.681 - 0.327 0.509 - 0.288 0.468 -
-${}^{{\dagger}}$ 0.640 0.778 - - - - - - - 0.318 0.496 - 0.261 0.430 -
-${}^{{\dagger}}$ 0.435 0.633 - 0.614 0.704 - 0.582 0.673 - 0.336 0.516 - 0.285 0.449 -
-${}^{{\dagger}}$ 0.461 0.660 - 0.637 0.723 - 0.605 0.693 - 0.374 0.549 - 0.334 0.493 -
+T5-base 0.410 0.608 - 0.600 0.682 - 0.556 0.641 - 0.260 0.450 - 0.238 0.418 -
+T5-large 0.548 0.713 - 0.631 0.716 - 0.595 0.681 - 0.327 0.509 - 0.288 0.468 -
+Temp-T5 0.640 0.778 - - - - - - - 0.318 0.496 - 0.261 0.430 -
+REMEMO-base 0.435 0.633 - 0.614 0.704 - 0.582 0.673 - 0.336 0.516 - 0.285 0.449 -
+REMEMO-large 0.461 0.660 - 0.637 0.723 - 0.605 0.693 - 0.374 0.549 - 0.334 0.493 -
 GPT-3.5 (ICL-SP) 0.598 0.699 - 0.640 0.668 - 0.512 0.506 - 0.303 0.409 - 0.365 0.453 -
 GPT-3.5 (ICL-CoT) 0.706 0.788 - 0.565 0.603 - 0.436 0.464 - 0.340 0.478 - 0.243 0.348 -
-$*{}^{*}$ (ICL-SP) 0.772 0.829 - 0.716 0.742 - 0.571 0.546 - 0.454 0.525 - 0.431 0.485 -
-$*{}^{*}$ (ICL-CoT) 0.821 0.865 - 0.662 0.693 - 0.618 0.636 - 0.388 0.480 - 0.352 0.447 -
+GPT-4 (ICL-SP) 0.772 0.829 - 0.716 0.742 - 0.571 0.546 - 0.454 0.525 - 0.431 0.485 -
+GPT-4 (ICL-CoT) 0.821 0.865 - 0.662 0.693 - 0.618 0.636 - 0.388 0.480 - 0.352 0.447 -
 Llama2-7B (ICL-SP) 0.415 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354 0.205 0.344 0.226 0.044 0.084 0.153
@@
 Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379 0.256 0.433 0.281 0.285 0.409 0.305
-Llama2-13B (SFT-TGSR) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649 0.424 0.522 0.432 0.356 0.469 0.399
+Llama2-13B (SFT-TGR) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649 0.424 0.522 0.432 0.356 0.469 0.399
 Table 4: Main results using different models and strategies.
 We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc).
-Note: (1) Results with ${}^{{\dagger}}$ are reported in the original papers.
-We fine-tune and evaluate the models on our dataset.
+Note: (1) Results with are reported in the original papers.
+We only fine-tune and evaluate the models on our dataset.
 (2) Results with * are evaluated on 1000 random test samples.
@@
 To obtain a fair comparison, we use Llama2-13B as the base model for all strategies.
-The basic strategy is in-context learning with standard Input/Output (ICL-SP), which is used to calculate the performance changes.
+The basic strategy used to calculate the performance changes is in-context learning with standard Input/Output prompt (ICL-SP).
 4.3 Main Results
-Can our strategy bring more reliable symbolic reasoning?
-Figure 4 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TGQA.
-It can be seen that LLMs learn better temporal reasoning with SFT than ICL.
-By providing CoTs with bootstrapping and our data augmentation strategies, the model performance gets further enhanced.
-Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies.
+Can our strategies bring more reliable reasoning over TGs?
+We show the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and graph data augmentation on TGQA (Figure 4).
+It can be seen that LLM learns TR better with SFT than ICL.
+By providing CoTs with bootstrapping and graph data augmentation strategies, the model performance gets further enhanced.
+Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies (Table 3).
 Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error).
-It can be seen in Table 3 that our strategies reduce all types of error rates (ER).
+It can be seen that our strategies reduce all types of error rate.
 Can our two-step framework lead to better temporal reasoning performance?
-Table 4 shows the comparison between different models on all datasets.
+We show the comparison between different models and strategies on all datasets (Table 4).
 First, we observed that among all the LLMs with ICL, GPT-4 has the strongest performance as expected.
 For the Llama2 family, larger models have better performance due to advanced context understanding and improved generalization.
-We also found that CoTs not always bring better temporal reasoning due to unreliable intermediate results and hallucinations.
-More importantly, our model (SFT-TGSR), which is based on Llama2-13B, shows a comparable or even better performance than GPT-4 on all datasets.
-To prove the effectiveness of our framework, we also provide the results of some alternative strategies, where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively.
-It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets.
-We hypothesize that this improvement in performance is because our two-step neuro-symbolic process provides an easier path toward answering the temporal questions for LLMs.
-Does these learned capabilities of temporal reasoning generalize to other tasks?
-Figure 5 shows the comparison between different strategies on TimeQA and TempReason.
-In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning.
-More importantly, we observe that per-training on TGQA improves the model performance compared with ICL.
-Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions.
-Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the temporal reasoning capability of LLMs.
+We also found that CoTs not always bring better TR due to unreliable intermediate results and hallucinations.
+From the results of some alternative strategies, where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and vanilla CoT distillation, respectively, we prove the effectiveness of our framework (SFT-TGR).
+More importantly, our model, which is based on Llama2-13B, shows a comparable or even better performance than GPT-4 on all datasets.
+It can be seen that the two-step framework brings a substantial performance improvement on different datasets.
+We hypothesize this performance improvement is because our two-step reasoning process provides an easier path toward answering the temporal questions for LLM.
+Do these learned capabilities of temporal reasoning generalize to other tasks?
+We show the comparison between different strategies (ICL with SP/CoT, SFT with TGQA/original data) on the two existing datasets, TimeQA and TempReason (Figure 5).
+Our framework learns the capabilities of text-to-TG translation and temporal graph reasoning, which brings better TR.
+More importantly, we observed that SFT on TGQA improves the model performance compared with ICL.
+It can be concluded that these necessary capabilities in TR are generalizable to different data distributions.
+Since TGQA is fully controllable and requires minimal supervision, we actually provide a general and effective way of TR capability improvement.
 4.4 Ablation Study
 We ablate different modules to see their contributions to the performance.
-Table 5 shows the performance comparison between different configurations.
+We show the performance comparison between different configurations (Table 5).
 To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
-From the ablation study, we obtain some insights: (1) LLMs can benefit from explicitly presented (temporal) graph which is intuitive, concise and structured.
-(2) Given a reliable graph, symbolic reasoning brings better performance than free-text CoTs.
+From the ablation study, we obtain some insights: (1) LLM can benefit from explicitly presented (temporal) graph which is intuitive, concise and structured.
+(2) Given a reliable graph, CoT bootstrapping with contrastive learning brings better performance than vanilla CoT distillation.
 (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning.
-(4) The introduction of external knowledge such as mathematics and commonsense can augment the generation.
+(4) The introduction of external knowledge such as mathematics and commonsense can further augment the generation.
 5 Related Work
-Language-based Temporal Reasoning  Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023).
+Language-based Temporal Reasoning.
+Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023; Vashishtha et al., 2020).
 The vision here is to help LMs understand temporal concepts.
@@
 Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
-Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023).
+Reasoning towards LMs.
+Although LLMs exhibit some emergent behaviors (Zhang et al., 2023a, 2024; Lai et al., 2024a, b; Wang et al., 2024; Chen et al., 2024a), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023).
 Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods.
@@
 On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
-Neuro-symbolic Reasoning Recent research has shown the advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Hamilton et al., 2022).
-In particular, knowledge graphs (KGs) as the foundation of (temporal) symbolic reasoning has been widely adopted in the past (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021).
-Moreover,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs.
-In this paper, we build the connection between language-based and temporal KG-based temporal symbolic reasoning.
+Reasoning over Knowledge Graphs.
+Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Yang et al., 2017; Sun et al., 2018; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021; Huang et al., 2023; Chen et al., 2023, 2024b; Wan et al., 2024; Li et al., 2024).
+Moreover,  (Liu et al., 2022; Yang et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs.
+In this paper, we build the connection between language-based and symbolic-based temporal reasoning.
 This connection brings the potential for extending these KG-based methods to language-based tasks.
-Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain datasets, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to symbolic reasoning.
+Different from existing methods (Luo et al., 2023; Zhang et al., 2023b; Yuan et al., 2023), which limit their application to other tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to deliberate reasoning.
 Config. EM F1 Acc
@@
 TG 0.675 0.754 0.711
-TG + SR(bs) 0.723 0.797 0.736
-TG + SR(bs + aug) 0.782 0.838 0.795
-TG + EK + SR(bs) 0.753 0.803 0.776
-TG + EK + SR(bs + aug) 0.797 0.850 0.819
-Table 5: Ablation study results on TGQA.
-To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
-TG: temporal graph, SR: symbolic reasoning, EK: external knwoledge, bs: bootstrapping, aug: data augmentation.
+TG + CoT(bs) 0.723 0.797 0.736
+TG + CoT(bs + aug) 0.782 0.838 0.795
+TG + EK + CoT(bs) 0.753 0.803 0.776
+TG + EK + CoT(bs + aug) 0.797 0.850 0.819
+Table 5: Ablation study results on TGQA with Llama2-13B. TG: temporal graph, CoT: chain of thought, EK: external knwoledge, bs: bootstrapping, aug: graph data augmentation.
 6 Conclusion
-TG-LLM, a new framework for language models, has been proposed to improve their performance on temporal reasoning.
-Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning steps.
-Experiments indicate that TG-LLM achieves better performance than existing methods.
-An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning.
-Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks.
+TG-LLM, a novel framework for LMs, has been proposed to improve their performance on temporal reasoning.
+To produce reliable final answers, our framework equips LLMs with the temporal graph and intermediate reasoning steps.
+Extensive experiments indicate that TG-LLM achieves better performance than existing pipelines.
+An interesting direction for future work is to extend it to more complex applications such as inductive and abductive reasoning.
+Due to the graph structure and capability of deliberate reasoning, it is promising to improve the model performance on these tasks.
 Limitations
@@
 For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range.
-In this way, we can further improve the quality of generation.
+In this way, we can further improve the quality of generated intermediate steps.
 Ethics Statement
@@
 None of such content reflects the opinions of the authors.
-Appendix A TGQA Dataset Statistics
-In this section, we show statistics for our TGQA dataset.
-Figure 6 shows the number of facts distribution in training and test set, which determine the complexity of the task to some extent.
-Figure 7 shows the types of questions distribution in training and test set, where Q0: "Which event occurred first, <Event_A> or <Event_B>?", Q1: "Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in the chronological order?", Q2: "How long did the event <Event_A> last?", Q3: "", Q4: "How much time passed between the start of <Event_A> and the start of <Event_B>?", Q5: "What happened right before/after <Event_A> started?", Q6: "When did the <Event_A> occur?", Q7: "True or false: <Event_A> and <Event_B> happened at the same year?", Q8: "True or false: <Event_A> was still happening when <Event_B> started?".
-Note that we give eight categories of question in Table 2, since in a strict sense Q0 and Q1 can be considered as the same type.
+Appendix A Dataset Statistics of TGQA
+In this section, we provide statistics for our TGQA dataset.
+We obtain 400 samples for training, 100 for validation and another 100 for test, with about 30 QA pairs in a single sample.
+We show the distribution of the number of facts (in one sample) in training (with validation) and test set (Figure 6).
+It determines the complexity of the TR tasks to some extent.
+We also show the distribution of question types in training (with validation) and test set (Figure 7), where Q0: "Which event occurred first, <Event_A> or <Event_B>?", Q1: "Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in the chronological order?", Q2: "How long did the event <Event_A> last?", Q3: "", Q4: "How much time passed between the start of <Event_A> and the start of <Event_B>?", Q5: "What happened right before/after <Event_A> started?", Q6: "When did the <Event_A> occur?", Q7: "True or false: <Event_A> and <Event_B> happened at the same year?", Q8: "True or false: <Event_A> was still happening when <Event_B> started?".
+Note that we give eight question categories in Table 2, since in a strict sense Q0 and Q1 can be considered as the same type.
 Among all the question types, Q1 has the largest portion since it has multiple variants.
 Similarly, Q7 with two variants also has a larger portion.
-To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores.
-We also provide examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.
+To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as the final scores.
+Additionally, we show examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.
+Figure 6: Number of facts distribution in TGQA.
+Figure 7: Types of questions distribution in TGQA.
+Appendix B Experiment Details
+In this section, we present more experiment details for further research. We include graph data augmentation examples, model versions and evaluation tasks & metrics.
 Entity Name Mapping
@@
 $\cdots$ $\cdots$
-Table 6: We use a global mapping for entity names from GPT-3.5 to avoid corpus leakage.
+Table 6: We use a global mapping for entity names from GPT-3.5 to avoid data leakage.
 External Knowledge
@@
 2 < 6 < 23 < 35
-Table 7: We integrate the necessary mathematics in context as external knowledge for temporal reasoning.
-Appendix B Experiment Details
-In this section, we present more experiment details for further research. We include data augmentation examples, model versions and evaluation tasks & metrics.
+Table 7: We integrate the necessary mathematics and commonsense in context as external knowledge for TR.
 Temporal Graph:
@@
 Parker was married to John Thompson)?
-CoT (symbolic): The duration for each event can be
+CoT: The duration for each event can be
 calculated as follows:
@@
 married to James Brown)?
-CoT (symbolic): The duration for each event can be
+CoT: The duration for each event can be
 calculated as follows:
@@
 A: True.
-Table 8: We propose special data augmentation strategies for symbolic reasoning over (temporal) graphs.
-Data Augmentation Examples  We show with an example our data augmentation strategies in Table 8.
+Table 8: We propose several graph data augmentation strategies for reasoning over temporal knowledge graph. The original TG is shown in Table 1. We highlight the changed information.
+Graph Data Augmentation.
+We show with an example our graph data augmentation strategies in Table 8.
 For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on graph-based QAs.
-For TGs with entity names/times mapping, we also need to change with the corresponding entity names/times in Graph-based QAs to ensure the consistency.
-Model Versions  All the model versions of LLMs used in our experiments are listed below:
-  • Llama2-7B (meta-llama/Llama-2-7b-hf)
-  • Llama2-13B (meta-llama/Llama-2-13b-hf)
-  • Llama2-70B (meta-llama/Llama-2-70b-hf)
-  • GPT-3.5 (gpt-3.5-turbo)
-  • GPT-4 (gpt-4-1106-preview)
-For the Llama2 family, all the model weights can be downloaded from the platform of Hugging Face. For the GPT models, all the model weights can be accessed through the OpenAI APIs.
-Evaluation Tasks & Metrics  In addition to TGQA, we consider the existing datasets TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries.
-To be specific, TimeQA  (Chen et al., 2021) contains two difficulty levels (see Table 9).
-The easy-level split tends to be information extraction while the hard-level split involves understanding the relation between different temporal expressions.
-TempReason  (Tan et al., 2023a) contains three levels (L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation) and three settings (OBQA: open-book QA, CBQA: close-book QA, ReasonQA: facts-based QA) (see Table 10).
-We observe that some stories in TempReason are incomplete which partially leads to the low accuracy of LLMs.
-We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
-Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations.
+To contrust, for TGs with entity names/times mapping, we need to change with the corresponding entity names/times in graph-based QAs to ensure the consistency.
+Evaluation Tasks & Metrics.
+Besides TGQA, we consider in experiments the two existing datasets TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a).
+Specifically, TimeQA contains two difficulty levels (Table 9).
+The easy-level split tends to be the information extraction task while the hard-level split involves understanding the relation between different temporal expressions.
+On the other hand, TempReason contains three levels (L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation) and three settings (OBQA: open-book QA, CBQA: close-book QA, ReasonQA: facts-based QA) (Table 10).
+We observed that some stories in TempReason are incomplete which partially leads to the low accuracy of LLMs.
 TimeQA
@@
 Table 10: Example questions of three difficulty levels under the OBQA setting in TempReason. L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation.
-Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format.
-This property actually hurts their performance under the commonly used metrics such as EM and F1.
-To solve this problem, we select from a candidate set the answer with the lowest perplexity from LLMs as their prediction.
-For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them.
-Since there are multiple categories of questions in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
+We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
+F1 and EM are two basic metrics for span-based QA tasks.
+However, the free-form prediction of LLMs might hurt their performance under these generation-based metrics.
+To solve this problem, we introduce perplexity-based accuracy, i.e., selecting from a candidate set the final answer with the lowest perplexity.
+For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best result among them.
+Since there are multiple question categories in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
     Model TGQA
 Q0 Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Total
-${}^{\ddagger}$ (0-shot SP) 0.801 0.435 0.660 0.355 0.930 0.955 0.710 0.315 0.617 0.642
+GPT-3.5 (0-shot SP) 0.801 0.435 0.660 0.355 0.930 0.955 0.710 0.315 0.617 0.642
 GPT-3.5 (1-shot SP) 0.842 0.411 0.656 0.555 0.924 0.650 0.500 0.353 0.495 0.598
-${}^{\ddagger}$ (0-shot CoT) 0.930 0.565 0.680 0.665 0.920 0.945 0.695 0.395 0.596 0.710
+GPT-3.5 (0-shot CoT) 0.930 0.565 0.680 0.665 0.920 0.945 0.695 0.395 0.596 0.710
 GPT-3.5 (1-shot CoT) 0.884 0.457 0.783 0.791 0.921 0.832 0.755 0.357 0.574 0.706
-${}^{{\ddagger}*}$ (0-shot SP) 0.801 0.689 0.920 0.620 0.930 0.940 0.758 0.575 0.737 0.776
-$*{}^{*}$ (1-shot SP) 0.820 0.660 0.870 0.630 0.910 0.990 0.760 0.610 0.700 0.772
-${}^{{\ddagger}*}$ (0-shot CoT) 0.950 0.692 0.901 0.881 0.919 0.980 0.778 0.647 0.858 0.848
-$*{}^{*}$ (1-shot CoT) 0.930 0.758 0.903 0.869 0.892 0.920 0.758 0.667 0.680 0.821
+GPT-4‡∗ (0-shot SP) 0.801 0.689 0.920 0.620 0.930 0.940 0.758 0.575 0.737 0.776
+GPT-4 (1-shot SP) 0.820 0.660 0.870 0.630 0.910 0.990 0.760 0.610 0.700 0.772
+GPT-4‡∗ (0-shot CoT) 0.950 0.692 0.901 0.881 0.919 0.980 0.778 0.647 0.858 0.848
+GPT-4 (1-shot CoT) 0.930 0.758 0.903 0.869 0.892 0.920 0.758 0.667 0.680 0.821
 Llama2-7B (1-shot SP) 0.554 0.216 0.163 0.147 0.353 0.924 0.683 0.115 0.579 0.415
@@
 Llama2-13B (1-shot SFT - TG) 0.878 0.515 0.656 0.866 0.835 0.845 0.726 0.369 0.383 0.675
-Llama2-13B (1-shot SFT - TG + SR(bs)) 0.947 0.732 0.747 0.825 0.931 0.987 0.495 0.466 0.649 0.753
-Llama2-13B (1-shot SFT - TG + SR(bs + aug)) 0.931 0.710 0.729 0.860 0.927 0.977 0.627 0.534 0.739 0.782
-Llama2-13B (1-shot SFT - TG + EK + SR(bs + aug)) 0.944 0.775 0.729 0.849 0.924 0.980 0.783 0.506 0.681 0.797
+Llama2-13B (1-shot SFT - TG + CoT(bs)) 0.947 0.732 0.747 0.825 0.931 0.987 0.495 0.466 0.649 0.753
+Llama2-13B (1-shot SFT - TG + CoT(bs + aug)) 0.931 0.710 0.729 0.860 0.927 0.977 0.627 0.534 0.739 0.782
+Llama2-13B (1-shot SFT - TG + EK + CoT(bs + aug)) 0.944 0.775 0.729 0.849 0.924 0.980 0.783 0.506 0.681 0.797
 […]
-We explain the categories of errors in human evaluation on CoT generation here (see Table 3).
+Human evaluation on CoT generation found out four types of error in total (Table 3).
 T1 means that LLM uses wrong information such as wrong start/end time during reasoning.
-T2 means that LLM makes logical errors, e.g., after determining "Event A ends before Event B starts", considering the statement as true that "Event A was still happening when Event B starts".
-T3 means that LLM makes errors on external knowledge, e.g., claiming that "the date 1978 is after 1983".
-T4 means that there exists error in the extracted TG which leads to the wrong conclusion.
+T2 suggests that LLM makes logical errors, e.g., mentioning "Event A ends before Event B starts" in CoT but determining the statement "Event A was still happening when Event B starts" to be true.
+T3 denotes that LLM makes errors on external knowledge, e.g., claiming that "the date 1978 is after 1983".
+T4 indicates that there exists errors in the extracted TG that lead to the wrong conclusion.
+Candidate Answer Generation.
+We involve candidate answers in the calculation of Acc.
+For TempReason, the authors provide negative answers since they perform time-sensitive reinforcement learning.
+That is, they use the score of the correct answers and wrong answers from the language model as reward to further finetune the model parameter.
+For TimeQA, we generate the candidates in the following way.
+Given a story, there are multiple related questions which share the same subject/object entity and relation.
+The correct answer changes with the query time.
+For example, "What position did George Washington hold in 1777/1790/1799?" Answer: "Commander in Chief/Presidency/Chancellor".
+We collect the answers of these related questions as candidates.
+More generally, if there are no such related questions, we will use all the entities in the corresponding TG as candidates.
+Model Versions.
+The versions of the LLMs used in our experiments are listed below.
+For the Llama2 family, all the model weights can be downloaded from the platform of Hugging Face.
+For the GPT models, all the model weights can be accessed through the OpenAI APIs.
+  • Llama2-7B (meta-llama/Llama-2-7b-hf)
+  • Llama2-13B (meta-llama/Llama-2-13b-hf)
+  • Llama2-70B (meta-llama/Llama-2-70b-hf)
+  • GPT-3.5 (gpt-3.5-turbo)
+  • GPT-4 (gpt-4-1106-preview)
 Appendix C Example Prompts
 In this section, we show example prompts used in our framework.
-To be specific, Table 12 shows an example of the graph-based story generation in TGQA; Table 13 presents an example of the automatic story-temporal graph alignment verification in TGQA; Table 14 provides an example of the temporal info identification in TimeQA; Table 15 illustrates an example of the graph construction in TimeQA; Table 16 gives an example of the automatic temporal graph-QA alignment verification in TimeQA; Table 17 depicts an illustrative case of the CoTs bootstrapping in TGQA.
-Appendix D Fine-grained Results
-We provide fine-grained results on TGQA in Table 11.
+Specifically, Table 12 shows an example of the graph-based story generation in TGQA; Table 13 presents an example of the automatic story-temporal graph alignment verification in TGQA; Table 14 provides an example of the temporal info identification in TimeQA; Table 15 illustrates an example of the graph construction in TimeQA; Table 16 gives an example of the automatic temporal graph-QA alignment verification in TimeQA; Table 17 depicts an illustrative case of the CoTs bootstrapping in TGQA.
+Appendix D Fine-grained Results of TGQA
+We provide the fine-grained results of TGQA in Table 11.
 In TGQA, there are nine types of questions as explained in Appendix A.
-We show all the models with different strategies here.
-Zero-shot performance of LLMs is also considered to investigate the effect of in-context examples.
+We show all the models with different strategies.
+Zero-shot performance is also considered to investigate the effect of in-context examples.
 Note that we do not include zero-shot performance of the Llama2 family since it is not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results.
@@
 Note that our model achieves the best overall performance by improving all the categories.
-To be specific, symbolic reasoning brings better performance for most categories.
-Furthermore, data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.
+Specifically, deliberate reasoning with CoT bootstrapping brings better performance for most categories.
+Furthermore, graph data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.
 (Molly Adams was born in Seattle) starts at 1896
@@
 When did the event (Molly Adams was married to Liam Thomas Dawson) end?
-Table 13: Example of the automatic story-temporal graph alignment verification in TGQA.
+Table 13: Example of the story-temporal graph alignment verification in TGQA.
 Knox Cunningham
@@
 time points (3 April 1909, 1930s, 1939, 1942, 1943, 1945, 1947, 1949, 1954, 29 July 1976).
-Table 15: Example of the graph construction in TimeQA.
+Table 15: Example of the temporal graph construction in TimeQA.
 3 April 1909: Knox Cunningham was born.
@@
 Which position did Knox Cunningham hold from May 1955 to Apr 1956?
-Table 16: Example of the automatic temporal graph-QA alignment verification in TimeQA.
+Table 16: Example of the temporal graph-QA alignment verification in TimeQA.
 Example:
@@
 Answer: Let’s think step by step.
-Table 17: Example of the CoTs bootstrapping in TGQA.
+Table 17: Example of the CoT bootstrapping in TGQA.

### v4

@@
 Recent studies have introduced various methods to mitigate these limitations.
-Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate temporal logic.
+Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal concepts and intricate temporal logic.
 In this paper, we propose TG-LLM, a novel framework towards language-based TR.
-Instead of reasoning over the original context, we adopt a latent representation, temporal graph (TG) that facilitates the TR learning.
+Instead of reasoning over the original context, we adopt a latent representation, temporal graph (TG) that enhances the learning of TR.
 A synthetic dataset (TGQA), which is fully controllable and requires minimal supervision, is constructed for fine-tuning LLMs on this text-to-TG translation task.
 We confirmed in experiments that the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks.
-On top of that, we teach LLM to perform deliberate reasoning over the TGs via Chain of Thought (CoT) bootstrapping and graph data augmentation.
+On top of that, we teach LLM to perform deliberate reasoning over the TGs via Chain-of-Thought (CoT) bootstrapping and graph data augmentation.
 We observed that those strategies, which maintain a balance between usefulness and diversity, bring more reliable CoTs and final results than the vanilla CoT distillation.
-[^1] The code is available at https://github.com/xiongsiheng/TG-LLM.
+[^1] Code and data are available at https://github.com/xiongsiheng/TG-LLM.
 1 Introduction
@@
 It is not just about understanding basic concepts such as ordering or duration; it extends to more intricate aspects, e.g., task planning or causal relation discovery.
-Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have emerged with some reasoning capabilities  (Huang and Chang, 2023).
+Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a) have emerged with some reasoning capabilities  (Huang and Chang, 2023).
 However, there is observation that they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023), preventing their applications from solving complex real-world problems.
-In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
-Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and temporal graph reasoning.
+In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Vashishtha et al., 2020; Qin et al., 2021).
+Figure 1: Our framework (TG-LLM) performs temporal reasoning in two steps: 1) Text-to-Temporal Graph translation: generate (relevant) temporal graph given the context and keyword (extracted from questions); 2) Temporal Graph Reasoning: perform Chain-of-Thought reasoning over the temporal graph.
 Recent works mainly adopt general approaches to investigate and improve the TR capability of LLMs.
-For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2023a).
-Similarly,  (Wei et al., 2023b) designs several specific types of prompts as prompt tuning.
+For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2022).
+Similarly,  (Wei et al., 2023) designs several specific types of prompts as prompt tuning.
 (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT).
@@
 We demonstrate the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks.
-Given a reliable TG, the key challenges of teaching LLMs TR include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR?
-Prior works  (Lewis et al., 2021) show that explicitly introducing knowledge in context enhances the performance of LLMs.
-In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps).
+Given a reliable TG, the key challenges of teaching TR to LLMs include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR?
+Prior work  (Lewis et al., 2021) shows that explicitly introducing knowledge into context enhances the performance of LLMs.
+In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., temporal relation and time gap between the timestamps, and the relative order of the gaps).
 (2) How can one teach LLM to perform deliberate reasoning?
@@
 For (ii), there is no guarantee for the correctness of generated intermediate steps especially with insufficient training data  (Yang et al., 2023b).
-We adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning.
-We further improve the model performance with graph data augmentation which mitigates the data deficiency in TR tasks.
+In this paper, we adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning.
+We further improve the model performance with graph data augmentation, which mitigates the data deficiency in TR tasks.
 To be specific, our contributions are summarized as follows:
@@
 Extensive experiments prove that our novel approach results in superior performance compared to the baselines.
-  • We design two approaches including Chain of Thought bootstrapping and graph data augmentation to teach LLM to generate consistent and faithful CoTs, which brings better performance than the vanilla CoT distillation.
+  • We design two approaches including Chain-of-Thought bootstrapping and graph data augmentation to teach LLM to generate consistent and faithful CoTs, which brings better performance than the vanilla CoT distillation.
   • We present a pipeline to create a synthetic dataset (TGQA) for question answering that requires TR.
@@
 We show in experiments that fine-tuning on our dataset benefits LLM on other TR tasks and benchmarks.
-Temporal Graph (sub; rel; obj; start/end; time):
+Temporal Graph [sub; rel; obj; start/end; time]:
 […]
@@
 $\cdots$
-Table 1: Each sample from TGQA dataset is in the form of (story, temporal graph, questions, answers).
+Table 1: Each sample from TGQA dataset is in the form of (temporal graph, story, questions, answers).
 Reasoning Type Question Answer
@@
 In this section, we present the construction pipeline for TGQA dataset that is fully controllable and requires minimal supervision for text-temporal graph alignment.
-Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of TR questions (Table 2), and the ground-truth timeline.
+Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have the ground-truth timelines and more diverse categories and types of TR questions (Table 2).
+More importantly, the pipeline can be used for various scenarios and tasks.
 We first split the large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), into subgraphs with a restriction on the number of events, and anonymize the entities to avoid data leakage.
@@
 Specifically, given a certain entity, we find its neighbors within three hops, and extract all the events happening between them.
-Since LLMs are supposed to do reasoning instead of memorization, we ensure that no overlapping exists between the events in training, validation and test sets.
+Since we hope LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the events in training, validation and test sets.
 Additionally, we notice the data leakage problem of LLMs, i.e., prior knowledge of the test data has been implicitly obtained from the extensive pre-training.
 Thus, an anonymization strategy, i.e., changing entity names into random ones of the same type, is adopted.
-We generate a global mapping for entity names with GPT-3.5.
+For each relation, we generate a global mapping of entity names with GPT-3.5.
 To avoid confusion, we adopt obscure names that do not exist in YAGO11k.
@@
 Based on the given subgraph, we generate a story and multiple QAs.
-We first ask GPT-3.5 to write a story based on the subgraph with the requirements to include all the events.
+We first ask GPT-3.5 to write a story based on the subgraph with the requirement to include all the events.
 It is observed that GPT-3.5 tends to ignore the end time of some events in the created story.
-To solve this problem, we separate the start and end time of the same event in prompt.
+To solve this problem, we separate the start and end time of the same event in the prompt.
 On the other hand, to obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction.
@@
 Step 3: Quality Control.
-In TGQA, we might introduce noise from the misalignment between the given subgraph and generated story.
-To reduce the noise, we propose a semi-automatic verification method.
-Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors, and only manually inspect the set of unanswered questions by LLM.
+In TGQA, noise might be introduced from the misalignment between the given subgraph and LLM-generated story.
+To address this problem, we propose a semi-automatic verification method.
+Fully manual inspection is expensive, but by first utilizing LLM we can narrow down potential errors, and only manually inspect the set of unanswered questions by LLM.
 Specifically, given a generated story, GPT-3.5 is queried on the time of each event in the graph.
-If it cannot give the correct answer, we consider the event as possibly missing in the story, which requires manual verification.
+If it cannot give the correct answer, we consider the event as possibly missing in the story, which requires further manual verification.
 We proved the effectiveness of our semi-automatic pipeline with fully manual verification of the test stories.
@@
 We provide these information to LLM for better alignment in TG construction, and verify the generated TG in a semi-automatic way.
-(1) Entity & Relation Extraction: To obtain the entity and relation in pre-defined QAs, we consider two ways: parsing with rules or using GPT-3.5.
+(1) Entity & Relation Extraction: To obtain the entity and relation in pre-defined QAs, we consider two strategies: parsing with rules or using GPT-3.5.
 Rule-based parsing, if applicable, is more efficient and reliable, which is prioritized in our experiments.
@@
 For those stories without explicit time expressions, we skip Step (2), and ask GPT-3.5 to provide the temporal order of events.
-Similar to TGQA, we verify the generated TG by first asking GPT-3.5 the pre-defined QAs, and then manually checking if GPT-3.5 fails.
-Figure 2: In Chain of Thoughts (CoTs) bootstrapping, we accept CoTs leading to correct final answers and sample them according to their contrastive learning scores.
+Similar to TGQA, we verify the generated TG by first asking GPT-3.5 the pre-defined QAs, and then manually checking the alignment if GPT-3.5 fails.
+Figure 2: In Chain-of-Thought (CoT) bootstrapping, we only accept CoTs that lead to correct final answers and sample them according to their contrastive learning scores to balance usefulness and diversity.
 Supervised Fine-tuning.
@@
 With these steps, the TG in context is transferred into a timeline with the alignment between entities, relations and times.
-We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021).
+We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023b) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021).
 The input and output of the model are the story and aligned timeline, respectively.
@@
 Specifically, we first prepare ICL examples for high-quality CoT generation.
-To facilitate learning, the pre-defined QAs are classified into different categories $\{Q_{1},Q_{2},\cdots$, $Q_{M}\}$.
+To facilitate the learning, the pre-defined QAs are classified into different categories $\{Q_{1},Q_{2},\cdots$, $Q_{M}\}$.
 For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$, where $g$, $e$, $q$, $a$ denote TG, external knowledge, question and answer, respectively, and ask both GPT-3.5 and GPT-4 to provide diverse CoTs $\{c_{j}\}^{N_{i}}_{j=1}$.
@@
 1).
-The $score$ design, inspired by  (Wang et al., 2023), considers both perplexity $P(\cdot)$ of the correct answer and plausibility growth $G(\cdot)$ (Eq.
+The $score$ design, inspired by  (Wang et al., 2023), considers both normalized probability $P(\cdot)$ of the correct answer and plausibility growth $G(\cdot)$ (Eq.
 2).
@@
 $P_{sample}(c_{k})=\operatorname{softmax}(score(c_{k}))$ (1)
-$score(c_{k})=\log P\left(a^{*}\mid g,e,q,c_{k}\right)+\gamma G\left(c_{k}\right)$ (2)
-$G\left(c_{k}\right)=\log\frac{P\left(a^{*}\mid g,e,q,c_{k}\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c_{k}\right)}$ (3)
-$P\left(a\mid g,e,q,c_{k}\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c_{k},t_{<l}\right)$ (4)
-where $c_{k}$ denotes the current CoT in consideration; $a^{\prime}$ denotes a certain wrong answer from the candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $t_{l}$ denotes the $l$-th token in $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
+$score(c_{k})=\log P\left(a^{*}\mid q^{\dagger},c_{k}\right)+\gamma G\left(c_{k}\right)$ (2)
+$G\left(c_{k}\right)=\log\frac{P\left(a^{*}\mid q^{\dagger},c_{k}\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid q^{\dagger},c_{k}\right)}$ (3)
+$\log P\left(a\mid q^{\dagger},c_{k}\right)=\frac{1}{|a|}\sum_{l}^{|a|}\log P\left(t_{l}\mid q^{\dagger},t_{<l}\right)$ (4)
+where $c_{k}$ denotes the current CoT in consideration; $q^{\dagger}:=\{g,e,q\}$; $a^{\prime}$ denotes a certain wrong answer from the candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $t_{l}$ denotes the $l$-th token in $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
+Figure 3: We further boost the model performance with several graph data augmentation strategies: remove irrelevant edges, use relation synonyms and change entities/times.
 3.2.2 Graph Data Augmentation
@@
 To address this issue, we propose several graph data augmentation strategies (Figure 3).
-Our framework involves two steps: text-to-TG translation and temporal knowledge graph reasoning.
-Notably, for the reasoning part, we train with ground-truth/verified TGs but infer on the estimated graphs.
-This discrepancy hurts the robustness of reasoning.
-Thus, we introduce some disturbances on the TGs during training.
+Our framework involves two steps: text-to-TG translation and temporal graph reasoning.
+Notably, for the reasoning part, the model is trained over ground-truth/verified TGs but infers on the estimated graphs.
+This discrepancy actually hurts the robustness of reasoning.
+Thus, we introduce some disturbances on the TG during training.
 Note that the type of disturbances should be carefully designed in order to avoid confusing the LLM.
@@
 $F^{\prime}\in\{f_{R}(F):P(c|f_{R}(g),e,q)\approx P_{0}\}_{F\in g}$ (6)
-Figure 3: We have several strategies for graph data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
 where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation names $f_{R}(\cdot)$.
@@
 4.1 Experimental Setup
-We demonstrate TG-LLM is a general framework by applying it to, besides TGQA, the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries.
+We demonstrate TG-LLM is a general framework by applying it to, besides TGQA, the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, and summaries.
 Examples from other datasets are listed in Appendix B.
@@
 Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluation, i.e., selecting from a candidate set the final answer with the lowest perplexity as the prediction.
-The construction of the candidate set for all datasets and detailed explanations are listed in Appendix B.
+The rationale and detailed construction of the candidate set for all datasets are listed in Appendix B.
 We primarily compare our framework with the leading LLMs, i.e., Llama2 Touvron et al.
-(2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023).
+(2023b), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023).
 Specifically, we ran the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and used APIs provided by OpenAI for GPT models.
@@
 Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
-4.2 Implementation Details
-We use Llama2-13B as the baseline due to limited computational resources.
-We inject two adapters with selectors into the base model for the text-to-TG translation and temporal knowledge graph reasoning.
-The adapters are trained in parallel.
-For inference, we first translate the original story into a temporal graph, and then perform reasoning on it, i.e., the adapters are used in sequence.
-For data generation, we use GPT-3.5 for story, TG and CoT generation, and the verification of stories and TGs.
-We use GPT-4 to create the ICL demonstrations of CoT generation, due to its high generation quality.
-All the prompt templates are given in Appendix C.
 Figure 4: Performance comparison between different CoT generation strategies on TGQA.
@@
 The basic strategy used to calculate the performance changes is in-context learning with standard Input/Output prompt (ICL-SP).
+4.2 Implementation Details
+We use Llama2-13B as the baseline due to limited computational resources.
+We inject two adapters with selectors into the base model for the text-to-TG translation and temporal graph reasoning.
+The adapters are trained in parallel.
+For inference, we first translate the original story into a temporal graph, and then perform reasoning on it, i.e., the adapters are used in sequence.
+For data generation, we use GPT-3.5 for story, TG and CoT generation, and the verification of stories and TGs.
+We use GPT-4 to create the ICL demonstrations of CoT generation, due to its high generation quality.
+All the prompt templates are given in Appendix C.
 4.3 Main Results
@@
 Language-based Temporal Reasoning.
-Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023; Vashishtha et al., 2020).
-The vision here is to help LMs understand temporal concepts.
-Existing methods mainly solve this problem with time-aware modeling.
-For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust TR.
-(Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge.
-Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
-Reasoning towards LMs.
-Although LLMs exhibit some emergent behaviors (Zhang et al., 2023a, 2024; Lai et al., 2024a, b; Wang et al., 2024; Chen et al., 2024a), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023).
-Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods.
-For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.
-(Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing final answers.
-On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
-Reasoning over Knowledge Graphs.
-Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Yang et al., 2017; Sun et al., 2018; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021; Huang et al., 2023; Chen et al., 2023, 2024b; Wan et al., 2024; Li et al., 2024).
-Moreover,  (Liu et al., 2022; Yang et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs.
-In this paper, we build the connection between language-based and symbolic-based temporal reasoning.
-This connection brings the potential for extending these KG-based methods to language-based tasks.
-Different from existing methods (Luo et al., 2023; Zhang et al., 2023b; Yuan et al., 2023), which limit their application to other tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to deliberate reasoning.
+Recently, language-based TR has gained substantial research focus  (Liu et al., 2023, 2024b; Chen et al., 2024a, b; Wang et al., 2024; Cheng et al., 2024a; Xia et al., 2024).
+The vision here is to help LMs understand temporal concepts and logic such that they can perform more complicate tasks.
+Existing methods mainly solve this problem with time-aware language modeling.
+For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose specific pre-training/fine-tuning strategies for robust TR.
+On the other hand,  (Ning et al., 2019; Zhou et al., 2020, 2021; Yang et al., 2020, 2023a) design auxiliary objectives to introduce external temporal knowledge.
+Although these methods made some progress, representation learning for the underlying structure and logic of TR are either ignored or not explicitly involved.
 Config. EM F1 Acc
@@
 Table 5: Ablation study results on TGQA with Llama2-13B. TG: temporal graph, CoT: chain of thought, EK: external knwoledge, bs: bootstrapping, aug: graph data augmentation.
+Reasoning towards LMs.
+Although LLM has exhibited some emergent behaviors (Zhang et al., 2024a; Lai et al., 2024a, b; Lyu et al., 2024), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is.
+Existing methods that try to elicit or enhance reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods.
+For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use in-context learning where some demonstrations including intermediate reasoning steps are provided in prompt.
+(Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with the intermediate thinking process before producing final answers.
+On the other hand,  (Schick et al., 2023; Kynoch et al., 2023; Xu et al., 2023) combine LMs with domain-specific external tools, empowering the model to perform more complex tasks that require reasoning and interactions with environment.
+Reasoning over Knowledge Graphs.
+Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Sun et al., 2018; Yang and Song, 2020; Huang et al., 2023; Wan et al., 2024; Li et al., 2024).
+Related work includes symbolic reasoning over temporal KGs  (Yang et al., 2022; Xiong et al., 2023, 2024), and language-based reasoning over static KGs  (Cheng et al., 2024b; Zhang et al., 2024b; Liu et al., 2024a; Zhao et al., 2024; Xu et al., 2024; Wei et al., 2024).
+In this paper, we build the connection between language-based and symbolic-based TR.
+This connection brings the potential for extending these KG-based methods to language-based tasks.
+Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to graph reasoning.
 6 Conclusion
@@
 An interesting direction for future work is to extend it to more complex applications such as inductive and abductive reasoning.
-Due to the graph structure and capability of deliberate reasoning, it is promising to improve the model performance on these tasks.
+Due to the graph structure and capability of deliberate reasoning, it is promising to improve the model performance on these tasks as well.
 Limitations
-Graph-augmented approaches such as TG-LLM help language models better learn related concepts from the perspective of graph.
+Graph-augmented approaches  (Jin et al., 2024; Fan et al., 2024; Shang and Huang, 2024; He and Hooi, 2024) including TG-LLM help language models better learn related concepts from the perspective of graph.
 Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021).
@@
 Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data.
-There can be opportunities such as simulating an environment to provide feedback to LLMs  (Yao et al., 2023).
+There can be opportunities such as simulating an environment to provide feedback to LLMs  (Hao et al., 2023).
 For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range.
-In this way, we can further improve the quality of generated intermediate steps.
+In this way, we can further improve the performance.
 Ethics Statement
@@
 None of such content reflects the opinions of the authors.
+Acknowledgements
+This work was supported by a sponsored research award by Cisco Research.
 Appendix A Dataset Statistics of TGQA
@@
 Additionally, we show examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.
-Figure 6: Number of facts distribution in TGQA.
-Figure 7: Types of questions distribution in TGQA.
-Appendix B Experiment Details
-In this section, we present more experiment details for further research. We include graph data augmentation examples, model versions and evaluation tasks & metrics.
 Entity Name Mapping
@@
 Table 7: We integrate the necessary mathematics and commonsense in context as external knowledge for TR.
+Appendix B Experiment Details
+In this section, we present more experiment details for further research. We include graph data augmentation examples, model versions and evaluation tasks & metrics.
 Temporal Graph:
@@
 For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on graph-based QAs.
-To contrust, for TGs with entity names/times mapping, we need to change with the corresponding entity names/times in graph-based QAs to ensure the consistency.
+To contrast, for TGs with entity names/times mapping, we need to change with the corresponding entity names/times in graph-based QAs to ensure the consistency.
 Evaluation Tasks & Metrics.
@@
 We observed that some stories in TempReason are incomplete which partially leads to the low accuracy of LLMs.
+We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
+F1 and EM are two basic metrics for span-based QA tasks.
+However, the free-form prediction of LLMs might hurt their performance under these generation-based metrics.
+To solve this problem, we introduce perplexity-based accuracy, i.e., selecting from a candidate set the final answer with the lowest perplexity.
+For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best result among them.
+Since there are multiple question categories in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
 TimeQA
@@
 Table 10: Example questions of three difficulty levels under the OBQA setting in TempReason. L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation.
-We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc).
-F1 and EM are two basic metrics for span-based QA tasks.
-However, the free-form prediction of LLMs might hurt their performance under these generation-based metrics.
-To solve this problem, we introduce perplexity-based accuracy, i.e., selecting from a candidate set the final answer with the lowest perplexity.
-For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best result among them.
-Since there are multiple question categories in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
     Model TGQA
@@
 Zero-shot performance is also considered to investigate the effect of in-context examples.
-Note that we do not include zero-shot performance of the Llama2 family since it is not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results.
+Note that we do not include zero-shot performance of the Llama2 family since they are not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results.
 For zero-shot learning results, the format of generation is not guaranteed.
 The original accuracy is very low since rule-based parser cannot handle it.
-Instead, we use GPT-3.5 as parser to extract the answer from the generation.
-Also, we only adopt one example in context due to context length limitation.
-It can be seen that in-context examples are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4.
+Instead, we use GPT-3.5 as parser to extract the answer from the generation (might introduce errors).
+Also, we only adopt one example in context due to context length limitation (long-context LLM performance could increase with more in-context demonstrations).
+It can be seen that in-context demonstrations are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4.
 Undoubtedly, GPT-4 obtains the best in-context learning performance on most categories.
 However, for some categories, the best performance of GPT-4 is around 0.6 or 0.7, which is not sufficiently well.
-On the other hand, we show that our strategies improve the performance step by step.
-Note that our model achieves the best overall performance by improving all the categories.
+On the other hand, we show that our strategies improve the performance in a progressive manner.
+Note that our model achieves the best overall performance by improving on all the categories.
 Specifically, deliberate reasoning with CoT bootstrapping brings better performance for most categories.